### v2

# Cuipo Jiang and Jingtian Song

Abstract
In this paper we first prove that the maximal ideal of the universal affine vertex operator algebra $V^{k}(sl_{n})$ for $k=-n+\frac{n-1}{q}$ is generated by two singular vectors of conformal weight $3q$ if $n=3$, and by one singular vector of conformal weight $2q$ if $n\geqslant 4$. We next determine the associated varieties of the simple vertex operator algebras $L_{k}(sl_{3})$ for all the non-admissible levels $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$. The varieties of the associated simple affine $W$-algebras $W_{k}(sl_{3},f)$, for nilpotent elements $f$ of $sl_{3}$, are also determined.
1 Introduction
Let $V$ be a vertex algebra. Recall from [Zhu96] that the Zhu’s $C_{2}$-algebra of $V$ is by definition the quotient space $R_{V}=V/C_{2}(V)$ of $V$, where $C_{2}(V)=\operatorname{span}_{{\mathbb{C}}}\{a_{-2}b\mid a,b\in V\}$. Then $R_{V}$ has the Poisson algebra structure given by
$\bar{a}\cdot\bar{b}=\overline{a_{-1}b},\qquad\{\bar{a},\bar{b}\}=\overline{a_{0}b},$
for $a,b\in V$, where $\bar{a}:=a+C_{2}(V)$ is the image of $a$ in $R_{V}$. The associated variety $X_{V}$ of $V$ is the reduced scheme $X_{V}=\operatorname{Specm}(R_{V})$ of $R_{V}$ [Ara12]. This fundamental invariant of $V$ has been extensively studied in [BFM91, Zhu96, ABD04, Miy04, Ara12, Ara15a, Ara15b, AM18a, AK18, AM17, AJM21, ADF+24, AFK24].
Let $\mathfrak{g}$ be a complex simple finite-dimensional Lie algebra and $V^{k}(\mathfrak{g})$ the universal affine vertex algebra at level $k\in{\mathbb{C}}$. By PBW Theorem, $R_{V^{k}(\mathfrak{g})}$ is isomorphic to the polynomial algebra ${\mathbb{C}}[\mathfrak{g}^{*}]$ over $\mathfrak{g}^{*}$. Thus the variety $X_{V}$ is the affine space $\mathfrak{g}^{*}$ with Kirillov-Kostant Poisson structure. For a graded quotient $V$ of $V^{k}(\mathfrak{g})$, the variety $X_{V}$ is a Poisson subscheme of $\mathfrak{g}^{*}$ which is $G$-invariant and conic, where $G$ is the adjoint group of $\mathfrak{g}$.
Denote by $L_{k}(\mathfrak{g})$ the simple quotient of $V^{k}(\mathfrak{g})$. Let $h$ be the Coxeter number of $\mathfrak{g}$, $h^{\vee}$ the dual Coxeter number, and $r^{\vee}$ the lacing number of $\mathfrak{g}$. It has been proved in [AJM21] that $L_{k}(\mathfrak{g})=V^{k}(\mathfrak{g})$ if and only if $X_{L_{k}(\mathfrak{g})}=\mathfrak{g}^{*}$. In [GK07], Gorelik and Kac proved that $V^{k}(\mathfrak{g})$ is not simple if and only if
$r^{\vee}(k+h^{\vee})\in{\mathbb{Q}}_{\geqslant 0}\backslash\left\{\frac{1}{m}\mid m\in{\mathbb{Z}}_{\geqslant 1}\right\}.$ (1.1)
Therefore,
[…]
It is known [Zhu96, DM06, Ara12] that $X_{L_{k}(\mathfrak{g})}=\{0\}$ if and only if $k$ is a non-negative integer. For an admissible $L_{k}(\mathfrak{g})$ [KW89], or equivalently, for an admissible number $k$ which by definition satisfies
$k+h^{\vee}=\frac{p}{q},\quad p,q\in{\mathbb{Z}}_{\geqslant 1},\ (p,q)=1,\ p\geqslant\begin{cases}h^{\vee}&\text{if }(r^{\vee},q)=1,\\ h&\text{if }(r^{\vee},q)\neq 1,\end{cases}$
the variety $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit in $\mathfrak{g}$. In particular, if $k$ is a non-degenerate admissible number, that is,
$q\geqslant\begin{cases}h&\text{if }(r^{\vee},q)=1,\\ r^{\vee}h^{\vee}&\text{if }(r^{\vee},q)=r^{\vee},\end{cases}$
then $X_{L_{k}(\mathfrak{g})}$ is the nilpotent cone ${\mathcal{N}}(\mathfrak{g})$ [Ara15a]. If $k=-h^{\vee}$, it follows from [FF92, FG04] that $X_{L_{k}(\mathfrak{g})}$ is also the nilpotent cone. It was observed in [AM18a, AM18b, AM17, ADF+24] that there are cases when $L_{k}(\mathfrak{g})$ is non-admissible and $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit, which provide examples for the conjecture in physics [XY21] that, in view of the 4D/2D duality, there should be a large list of non-admissible simple affine vertex algebras whose associated varieties are the closures of some nilpotent orbits. There are also cases where $X_{L_{k}(\mathfrak{g})}$ is neither $\mathfrak{g}^{*}$ nor contained in the nilpotent cone ${\mathcal{N}}(\mathfrak{g})$. It was proved in [AM17] that $X_{L_{-1}(sl_{n})}$, $n\geqslant 4$, $X_{L_{-m}(sl_{2m})}$, $m\geqslant 2$, $X_{L_{2-r}(so_{2r})}$, $r\in 2{\mathbb{Z}}_{+}+1$, are closures of sheets. In [AFK24], it was shown that the associated variety of a simple affine vertex algebras is contained in the closure of the Dixmier sheet when a chiralization of generalized Grothendieck’s simultaneous resolution exists. The variety $X_{L_{-\frac{5}{2}}}(sl_{4})$ of $L_{-\frac{5}{2}}(sl_{4})$ was given in [Fas22], which turns out to be the closure of a Jordan class of $sl_{4}$.
In general, the problem of determining $X_{L_{k}(\mathfrak{g})}$ is widely open. Up to now, the variety $X_{L_{k}(\mathfrak{g})}$ has been completely determined for all $k\in{\mathbb{C}}$ only when $\mathfrak{g}=sl_{2}$. In this paper, we will determine $X_{L_{k}(sl_{3})}$ for all non-admissible numbers $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$. Then together with the known results $X_{L_{k}(\mathfrak{g})}$ for the critical case and admissible cases given in [FF92, FG04, Ara12], we will obtain a complete characterization of $X_{L_{k}(sl_{3})}$ for all $k\in{\mathbb{C}}$.
Another question we are concerned about in this paper is the characterization of the maximal ideal of $V^{k}(sl_{n})$, where $k=-n+\frac{n-1}{q}$, $n\geqslant 3$, $q\geqslant 2$, $(n-1,q)=1$. If $q=1$, that is, $k=-1$, the maximal ideal of $V^{-1}(sl_{3})$ is generated by two singular vectors of conformal weight 3 [AP08, AKMF+20], and the maximal ideal of $V^{-1}(sl_{n})$ is generated by one singular vector of conformal weight 2 [AP08, AM17]. By using the character formulas given in [KT00], the fact that the coefficients as values of Kazhan-Lusztig (or inverse KL) polynomials at 1 in the formulas depend only on the corresponding Coxeter group ( see also Theorem 11 in [Fie06] from the category point of view), and the results for the $k=-1$ case in [KW18, BKK24, AP08, AM17, AKMF+20], we prove our first main result as follows (also see Theorem 3.8).
Theorem 1.1.
Let $k=-n+\frac{n-1}{q}$ with $n\geqslant 3$, $q\geqslant 2$, $(n-1,q)=1$. Then
$\widehat{R}{\rm ch}L_{k}(sl_{n})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{q\gamma}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
Furthermore, we have
  1. (1)
    If $k=-3+\frac{2}{2m+1}$ with $m\geqslant 1$, then the maximal ideal of $V^{k}(sl_{3})$ is generated by two singular vectors $v^{1}$ and $v^{2}$ of weights $k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2}$ and $k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2}$, respectively.
  2. (2)
    If $k=-n+\frac{n-1}{q}$ with $n\geqslant 4$, $q\geqslant 2$, and $(q,n-1)=1$, then the maximal ideal of $V^{k}(sl_{n})$ is generated by one singular vector $w^{1}$ of weight $k\Lambda_{0}-2q\delta+\alpha_{1}+2\alpha_{2}+\cdots+2\alpha_{n-2}+\alpha_{n-1}$.
Let $\theta$ be the highest root of $\mathfrak{g}$. Our second main result is (also see Theorem 4.2):
Theorem 1.2.
The variety $X_{L_{-1}(sl_{3})}$ is the closure of the Dixmier sheet ${\mathbb{S}}_{\mathfrak{l}_{1}}=G.{\mathbb{C}}^{*}\lambda$, i.e.,
$X_{L_{-1}(sl_{3})}=\overline{{\mathbb{S}}_{\mathfrak{l}_{1}}}=\overline{G.{\mathbb{C}}^{*}\lambda}=G.{\mathbb{C}}^{*}\lambda\cup\overline{{\mathcal{O}}_{\min}},$
where ${\mathfrak{l}}_{1}=\mathfrak{h}\oplus{\mathbb{C}}e_{\theta}\oplus{\mathbb{C}}f_{\theta}$, $\lambda=h_{1}-h_{2}$, and ${\mathcal{O}}_{\min}$ is the minimal non-zero nilpotent orbit of $sl_{3}({\mathbb{C}})$. In particular, $X_{L_{-1}(sl_{3})}$ is irreducible and $\dim X_{L_{-1}(sl_{3})}=5$.
For $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, we have our third main result as follows (see also Theorem 4.9).
Theorem 1.3.
For $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, the variety
$X_{L_{k}(sl_{3})}=\overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
The proof of Theorem 1.3 is much more complicated than the $k=-1$ case. The main difficulty is that although we could prove that the maximal ideal of $L_{k}(sl_{3})$ is generated by two singular vectors $v^{1}$ and $v^{2}$ of conformal weight $3(2m+1)$ by Theorem 1.1, it is almost impossible to write these singular vectors out as sums of PBW terms for general $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$. This means that we could not use their structure, as what we do with the $k=-1$ case, to obtain the zero locus of the graded ideal $\overline{{\mathcal{I}}_{k}}$ of the $C_{2}$-algebra $R_{V^{k}(sl_{3})}$ determined by these two singular vectors. Our main idea is to analyze coefficients of some critical leading terms. Then further prove that $X_{L_{k}(sl_{3})}=\overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}$. Our method here provides a possible way to determine varieties of a simple affine vertex operator algebra without knowing the exact structure of its singular vectors. We will study the associated varieties of the simple vertex operator algebras $V^{k}(sl_{n})$ for $n\geqslant 4$ and general non-admissible numbers $k$ in subsequent papers.
Representations of $L_{k}(sl_{3})$ have been extensively studied in [AP08, Ada16, AFR17, Cre18, FLLZ18, AM21, AK21, AKR21, FKR21, FR22, KRW22, AFP23a, CRR23, ACG24], etc. We here have the following conjecture for $L_{k}(sl_{3})$ (also see Conjecture 4.11).
Conjecture 1.4.
Let $k=-3+\frac{2}{2m+1}$ with $m\geqslant 0$. Then
$\{L(t\bar{\Lambda}_{1}-\frac{2i}{3}\bar{\Lambda}_{2}),L(t\bar{\Lambda}_{2}-\frac{2i}{3}\bar{\Lambda}_{1}),L(\bar{\Lambda}_{1}-(t+\frac{2i}{3}+1)\bar{\Lambda}_{2}),~{}t\in{\mathbb{C}},i=0,1,\cdots,2m\}$
provides a complete list of irreducible $A(L_{k}(sl_{3}))$-modules in the category ${\mathcal{O}}$.
When $m=0$, the conjecture is true by [AP08]. When $m=1$, we could verify the conjecture by computer programming.
Let $f$ be a nilpotent element of $\mathfrak{g}=sl_{3}$. Then $f$ is minimal or regular. For $k\in{\mathbb{C}}$, let $W^{k}(sl_{3},f)$ be the universal affine $W$-vertex algebra associated to $f$ [FF92, KRW03]. Denote by $W_{k}(sl_{3},f)$ the simple quotient of $W^{k}(sl_{3},f)$.
Let $f$ be a minimal nilpotent element of $\mathfrak{g}=sl_{3}$. If $k=-3+\frac{2m+1}{2}$, $m\geqslant 1$, it was shown in [Ara13] that $X_{W_{k}(sl_{3},f)}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational. If $k=-1$, it has been proved in [AKF+18] that $W_{-1}(sl_{3},f)$ is isomorphic to the rank one Heisenberg vertex operator algebra. So $X_{W_{-1}(sl_{3},f)}$ is one-dimensional. We have the following result for $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$ (also see Theorem 5.3).
Theorem 1.5.
Let $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, and $f$ a minimal nilpotent element of $sl_{3}$. Then
$X_{W_{k}(sl_{3},f)}=\left\{\left[\begin{array}[]{ccc}a&b&3(\mu^{2}-a^{2})\\ 0&-2a\mu&c\\ 1&0&a\end{array}\right]|~{}\mu,a,b,c\in{\mathbb{C}},\ bc=2(a-\mu)(2a+\mu)^{2}\right\}.$
In particular, $\dim X_{W_{k}(sl_{3},f)}=3$. Furthermore, $W_{k}(sl_{3},f)$ is not quasi-lisse.
Let $f$ be a regular nilpotent element of $\mathfrak{g}=sl_{3}$. If $k=-3+\frac{p}{q}$, $p,q\geqslant 3$, $(p,q)=1$, by [Ara15b], $X_{W_{k}(sl_{3})}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational. Our fifth main result is given as follows (also see Theorem 5.4).
Theorem 1.6.
Let $k+3=\frac{2}{2m+1}$ or $\frac{2m+1}{2}$, $m\geqslant 1$, and $f$ a regular nilpotent element of $sl_{3}$. Then
$X_{W_{k}(sl_{3})}=\left\{\left[\begin{array}[]{ccc}0&\frac{3}{4}\mu^{2}&\frac{1}{2}\mu^{3}\\ 2&0&\frac{3}{4}\mu^{2}\\ 0&2&0\end{array}\right]|~{}\mu\in{\mathbb{C}}\right\}.$
In particular, $\dim X_{W_{k}(sl_{3})}=1$. Furthermore, $W_{k}(sl_{3},f)$ is not quasi-lisse.
The rest of this paper is organized as follows. In Section 2, we recollect some concepts and results that will be needed later. In Section 3, we study singular vectors and maximal ideals of $V^{k}(sl_{n})$ for $k=-n+\frac{n-1}{q}$ with $n\geqslant 3$, $q\geqslant 1$, $(n-1,q)=1$. Section 4 is dedicated to determine the associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$. In Section 5, we obtain the associated varieties of the simple affine $W$-algebras $W_{k}(sl_{3},f)$.
Acknowledgements
We appreciate Tomoyuki Arakawa, Dražen Adamović, and Anne Moreau for valuable discussions and comments. In particular, the first author is grateful to Dražen Adamović for showing her the very helpful result of Theorem 8.2 in his paper [AFP23b] joint with Pierluigi Möseneder Frajria and Paolo Papi, and to Tomoyuki Arakawara for telling her the papers [Fie06] and [BKK24], and for the discussions.
2 Preliminaries
2.1 $C_{2}$-algebras and associated varieties
Let $V$ be a vertex algebra over ${\mathbb{C}}$ [FLM89, FHL93, LL04]. Recall that a subset $S$ of $V$ is called a strongly generating set of $V$ if $V$ is linearly spanned by elements of the form:
[…]
Furthermore, if $S$ is a finite set, then $V$ is called finitely strongly generated by $S$.
For a vertex algebra $V$, define $C_{2}(V)=\operatorname{span}_{{\mathbb{C}}}\{a_{-2}b\mid a,b\in V\}$ and
$R_{V}=V/C_{2}(V).$
Then $R_{V}$ is a Poisson algebra [Zhu96] with relations:
$\bar{a}\cdot\bar{b}=\overline{a_{-1}b},\qquad\{\bar{a},\bar{b}\}=\overline{a_{0}b},$
for $a,b\in V$, where $\bar{a}=a+C_{2}(V)$. It is easy to see that $V$ is finitely strongly generated if and only if $R_{V}$ is finitely generated. $V$ is called $C_{2}$-cofinite if $R_{V}$ is finite-dimensional [Zhu96].
Let $V$ be a finitely strongly generated vertex algebra. By definition [Ara12], the associated variety of $V$ is the reduced scheme
$X_{V}={\rm Specm}(R_{V}).$
Recall from [Ara12] that $V$ is called lisse if $X_{V}=\{0\}$. We have the following result.
Theorem 2.1 ([Ara12]).
Let $V$ be a finitely strongly generated vertex algebra. Then $V$ is lisse if and only if $V$ is $C_{2}$-cofinite.
Since $R_{V}$ is a finitely generated Poisson algebra, the variety $X_{V}$ is a union of symplectic leaves [BG03, AK18].
Definition 2.2 ([AK18]).
Let $V$ be a vertex algebra. Then $V$ is called quasi-lisse if $X_{V}$ has finitely many symplectic leaves.
Theorem 2.3 ([AK18]).
Let $V$ be a finitely strongly generated vertex algebra. If $V$ is quasi-lisse, then $V$ has finitely many ordinary modules.
2.2 Affine vertex algebras
Let $\mathfrak{g}$ be a finite-dimensional simple Lie algebra over ${\mathbb{C}}$ with the following normalized bilinear form:
$(\cdot|\cdot)=\displaystyle{\frac{1}{2h^{\vee}}\times}\text{ Killing form of }\mathfrak{g}.$
Let $\widehat{\mathfrak{g}}=\mathfrak{g}[t,t^{-1}]\oplus{\mathbb{C}}K$ be the associated affine Kac-Moody algebra with the commutation relations:
$[x\otimes t^{m},y\otimes t^{n}]=[x,y]\otimes t^{m+n}+m(x|y)\delta_{m+n,0}K,\quad[K,\widehat{\mathfrak{g}}]=0,$
where $x,y\in\mathfrak{g}$ and $m,n\in{\mathbb{Z}}$. For $x\in\mathfrak{g}$ and $m\in{\mathbb{Z}}$, we will write $x(m)$ for $x\otimes t^{m}$.
For $k\in{\mathbb{C}}$, set
$V^{k}(\mathfrak{g})=U(\widehat{\mathfrak{g}})\otimes_{U(\mathfrak{g}[t]\oplus{\mathbb{C}}K)}{\mathbb{C}}_{k},$
where ${\mathbb{C}}_{k}$ is the one-dimensional representation of $\mathfrak{g}[t]\oplus{\mathbb{C}}K$ on which $K$ acts as multiplication by $k$ and $\mathfrak{g}[t]$ acts trivially. By PBW Theorem, we have
$V^{k}(\mathfrak{g})\cong U(\mathfrak{g}\otimes t^{-1}{\mathbb{C}}[t^{-1}])=U(t^{-1}\mathfrak{g}[t^{-1}]).$ (2.1)
The space $V^{k}(\mathfrak{g})$ is naturally graded,
$V^{k}(\mathfrak{g})=\bigoplus_{n\in{\mathbb{Z}}_{\geqslant 0}}V^{k}(\mathfrak{g})_{n},$
where the grading is defined by
$\deg(x^{i_{1}}(-n_{1})\ldots x^{i_{r}}(-n_{r}){\bf 1})=\sum_{i=1}^{r}n_{i},\quad r\geqslant 0,\;x^{i_{j}}\in\mathfrak{g},$
with ${\bf 1}$ the image of $1\otimes 1$ in $V^{k}(\mathfrak{g})$. We have $V^{k}(\mathfrak{g})_{0}={\mathbb{C}}{\bf 1}$, and we may identify $\mathfrak{g}$ with $V^{k}(\mathfrak{g})_{1}$ via the linear isomorphism defined by $x\mapsto x(-1){\bf 1}$.
It is well-known that $V^{k}(\mathfrak{g})$ has a unique vertex algebra structure such that ${\bf 1}$ is the vacuum vector,
$x(z)=Y(x\otimes t^{-1},z)=\sum\limits_{n\in{\mathbb{Z}}}x(n)z^{-n-1},$
and
$[T,x(z)]=\partial_{z}x(z)$
for $x\in\mathfrak{g}$, where $T$ is the translation operator. Notice that $x(n)$ acts on $V^{k}(\mathfrak{g})$ by left multiplication. Then one can view $x(n)$ as an endomorphism of $V^{k}(\mathfrak{g})$. The vertex algebra $V^{k}(\mathfrak{g})$ is called the universal affine vertex algebra associated with $\mathfrak{g}$ at level $k$ [FZ92, Zhu96, LL04].
If $k+h^{\vee}\not=0$, the vertex algebra $V^{k}(\mathfrak{g})$ is a vertex operator algebra by the Sugawara construction. More specifically, set
$S=\displaystyle{\frac{1}{2}}\sum_{i=1}^{d}x_{i}(-1)x^{i}(-1){\bf 1},$
where $d=\dim\mathfrak{g}$, and $\{x_{i}\colon i=1,\ldots,d\}$ is the dual basis of a basis $\{x^{i}\colon i=1,\ldots,\dim\mathfrak{g}\}$ of $\mathfrak{g}$ with respect to the bilinear form (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ). Then the vector $\omega=\displaystyle{\frac{S}{k+h^{\vee}}}$ is a conformal vector of $V^{k}(\mathfrak{g})$ with central charge
$c(k)=\displaystyle{\frac{k\dim\mathfrak{g}}{k+h^{\vee}}}.$
Any graded quotient of $V^{k}(\mathfrak{g})$ as $\widehat{\mathfrak{g}}$-module has the structure of a quotient vertex algebra. In particular, the unique simple graded quotient $L_{k}(\mathfrak{g})$ is a vertex algebra, and is called the simple affine vertex algebra associated with $\mathfrak{g}$ at level $k$.
2.3 Zhu’s algebra of affine vertex algebras
For a ${\mathbb{Z}}_{\geqslant 0}$-graded vertex algebra $V=\oplus_{n\geqslant 0}V_{n}$, let $O(V)$ be the subspace of $V$ linearly spanned by elements
$u\circ v=\sum\limits_{j=0}^{\infty}\binom{n}{j}u_{j-2}v,$
for $u\in V_{n},n\in{\mathbb{Z}}_{\geqslant 0}$, $v\in V$. The Zhu’s algebra [Zhu96]
$A(V)=V/O(V)$
is a unital associative algebra with the multiplication induced from
$u*v=\sum\limits_{j=0}^{\infty}\binom{n}{j}u_{j-1}v,$
for $u\in V_{n},n\in{\mathbb{Z}}_{\geqslant 0}$, $v\in V$. It is known from [FZ92] that the Zhu’s algebra of $V^{k}(\mathfrak{g})$ is the universal enveloping algebra $U(\mathfrak{g})$ of $\mathfrak{g}$. Let ${\mathcal{I}}_{k}$ be the maximal ideal of $V^{k}(\mathfrak{g})$. Then
$L_{k}(\mathfrak{g})=V^{k}(\mathfrak{g})/{\mathcal{I}}_{k}.$
Denote by $I_{k}$ the image of ${\mathcal{I}}_{k}$ in $A(V^{k}(\mathfrak{g}))$, then
$A(L_{k}(\mathfrak{g}))=U(\mathfrak{g})/I_{k}.$
Let $\mathfrak{h}$ be the Cartan subalgebra of $\mathfrak{g}$, and $\mathfrak{g}=\mathfrak{n}^{+}+\mathfrak{h}+\mathfrak{n}^{-}$ be a fixed triangular decomposition of $\mathfrak{g}$. Set
$U(\mathfrak{g})^{\mathfrak{h}}=\{u\in U(\mathfrak{g})|[h,u]=0,\ {\rm for~{}all}\ h\in\mathfrak{h}\}$
and let
$\mathfrak{p}:U(\mathfrak{g})^{\mathfrak{h}}\to U(\mathfrak{h})$
be the Harish-Chandra projection map, which is the restriction of the projection map: $U(\mathfrak{g})=U(\mathfrak{h})\oplus(\mathfrak{n}^{-}U(\mathfrak{g})\oplus U(\mathfrak{g})\mathfrak{n}^{+})\to U(\mathfrak{h})$ to $U(\mathfrak{g})^{\mathfrak{h}}$. It is known that $\mathfrak{p}$ is an algebra homomorphism [Hum72]. For a two-sided ideal $I$ of $U(\mathfrak{g})$, the characteristic variety of $I$ is defined as [Jos06]
${\mathcal{V}}(I)=\{\lambda\in\mathfrak{h}^{*}|p(\lambda)=0\ {\rm for~{}all}\ p\in\mathfrak{p}(I\cap U(\mathfrak{g})^{\mathfrak{h}})\}.$
For $\lambda\in\mathfrak{h}^{*}$, let $L(\lambda)$ be the irreducible module of $V^{k}(\mathfrak{g})$ induced from the simple highest weight $\mathfrak{g}$-module $L(\bar{\lambda})$. Then by [Zhu96], $L(\lambda)$ is a module of $L_{k}(\mathfrak{g})$ if and only if $I_{k}(L(\bar{\lambda}))=0$. We have the following result from [Ara16].
Proposition 2.4 ([Ara16]).
$L(\lambda)$ is an $L_{k}(\mathfrak{g})$-module if and only if $\bar{\lambda}\in{\mathcal{V}}(I_{k})$.
2.4 Associate graded vertex Poisson algebras of affine vertex algebras
It is known by Li [Li05] that any vertex algebra $V$ admits a canonical filtration $F^{\bullet}V$, called the Li filtration of $V$.
For a quotient $V$ of $V^{k}(\mathfrak{g})$, $F^{\bullet}V$ is described as follows. The subspace $F^{p}V$ is spanned by the elements
$y_{1}(-n_{1}-1)\cdots y_{r}(-n_{r}-1)\mathbf{1}$
with $y_{i}\in\mathfrak{g}$, $n_{i}\in{\mathbb{Z}}_{\geqslant 0}$, $n_{1}+\cdots+n_{r}\geqslant p$. We have
$V=F^{0}V\supset F^{1}V\supset\cdots,\quad\bigcap_{p}F^{p}V=0,$
$TF^{p}V\subset F^{p+1}V,$
$a_{n}F^{q}V\subset F^{p+q-n-1}V~{}\text{for}~{}a\in F^{p}V,\ n\in{\mathbb{Z}},$
$a_{n}F^{q}V\subset F^{p+q-n}V~{}\text{for}~{}a\in F^{p}V,\ n\geqslant 0.$
Here we set $F^{p}V=V$ for $p<0$.
Let ${\rm gr}^{F}V=\bigoplus_{p}F^{p}V/F^{p+1}V$ be the associated graded vector space. The space ${\rm gr}^{F}V$ is a vertex Poisson algebra by
$\sigma_{p}(a)\sigma_{q}(b)=\sigma_{p+q}(a_{-1}b),$
$T\sigma_{p}(a)=\sigma_{p+1}(Ta),$
$\sigma_{p}(a)_{n}\sigma_{q}(b)=\sigma_{p+q-n}(a_{n}b)$
for $a,b\in V$, $n\geqslant 0$, where $\sigma_{p}\colon F^{p}(V)\rightarrow F^{p}V/F^{p+1}V$ is the principal symbol map. In particular, the space $\operatorname{gr}^{F}V$ is a $\mathfrak{g}[t]$-module by the correspondence
$\mathfrak{g}[t]\ni x(n)\longmapsto\sigma_{0}(x)\in{\rm End}(\operatorname{gr}^{F}V)$ (2.2)
for $x\in\mathfrak{g}$, $n\geqslant 0$.
The filtration $F^{\bullet}V$ is compatible with the grading: $F^{p}V=\bigoplus\limits_{n\in{\mathbb{Z}}_{\geqslant 0}}F^{p}V_{n}$, where $F^{p}V_{n}:=V_{n}\cap F^{p}V$.
2.5 Zhu’s $C_{2}$-algebras and associated varieties of affine vertex algebras
Let $F^{\bullet}V$ be given as above. We have [Li05, Lemma 2.9]
$F^{p}V={\rm span}_{\mathbb{C}}\{a_{-i-1}b\colon a\in V,i\geqslant 1,b\in F^{p-i}V\}$
for all $p\geqslant 1$. In particular,
$F^{1}V=C_{2}(V),\qquad R_{V}=V/C_{2}(V)=F^{0}V/F^{1}V\subset{\rm gr}^{F}V.$
It is easily seen that
$F^{1}V^{k}(\mathfrak{g})=C_{2}(V^{k}(\mathfrak{g}))=t^{-2}\mathfrak{g}[t^{-1}]V^{k}(\mathfrak{g}).$
The following map defines an isomorphism of Poisson algebras
$\begin{array}[t]{rcl}{\mathbb{C}}[\mathfrak{g}^{*}]\cong S(\mathfrak{g})&\longrightarrow&R_{V^{k}(\mathfrak{g})},\\[1.99997pt] \mathfrak{g}\ni x&\longmapsto&x(-1){\bf 1}+t^{-2}\mathfrak{g}[t^{-1}]V^{k}(\mathfrak{g}).\end{array}$
So
$R_{V^{k}(\mathfrak{g})}\cong{\mathbb{C}}[\mathfrak{g}^{*}]$
and
$X_{V^{k}(\mathfrak{g})}\cong\mathfrak{g}^{*}.$
More generally, let $V$ be a quotient of $V^{k}(\mathfrak{g})$ by an ideal $N$, then we have
$R_{V}\cong{\mathbb{C}}[\mathfrak{g}^{*}]/\overline{N}$ (2.3)
as Poisson algebras, where $\overline{N}$ is the image of $N$ in $R_{V^{k}(\mathfrak{g})}={\mathbb{C}}[\mathfrak{g}^{*}]$. Then $X_{V}$ is just the zero locus of $\overline{N}$ in $\mathfrak{g}^{*}$. It is a closed $G$-invariant conic subset of $\mathfrak{g}^{*}$. Identifying $\mathfrak{g}^{*}$ with $\mathfrak{g}$ via the bilinear form (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ), one may view $X_{V}$ as a subvariety of $\mathfrak{g}$.
2.6 Sheet and Jordan classes
In this subsection, we recall Jordan classes and sheets of a semisimple Lie algebras following [TY05]. Let $\mathfrak{g}$ be as above. Denote by ${\mathcal{S}}$ (resp. ${\mathcal{N}}$) the set of semisimple (resp. nilpotent) elements of ${\mathfrak{g}}$. For $x\in\mathfrak{g}$, set $\mathfrak{g}^{x}=\{y\in\mathfrak{g}|[x,y]=0\}$, and denote by $x_{s}$ and $x_{n}$ its semisimple and nilpotent components, respectively. For $x,y\in\mathfrak{g}$, we say that $x$ and $y$ are $G$-Jordan equivalent if there exists $\alpha\in G$ such that
$\mathfrak{g}^{y_{s}}=\mathfrak{g}^{\alpha(x_{s})}=\alpha(\mathfrak{g}^{x_{s}}),\qquad y_{n}=\alpha(x_{n}).$
This defines an equivalence relation on $\mathfrak{g}$. The equivalence class of $x$, denoted by $J_{G}(x)$, is called the Jordan class of $x$ in $\mathfrak{g}$. A Jordan class is clearly a $G$-stable set.
For any Lie subalgebra $\mathfrak{u}\subset\mathfrak{g}$, set
$\mathfrak{z}(\mathfrak{u})=\{x\in\mathfrak{u}|[x,y]=0,\ y\in\mathfrak{u}\}.$
Denote by $\mathfrak{u}^{reg}$ the set of elements $y\in\mathfrak{u}$ such that the dimension of $\mathfrak{g}^{y}$ is minimal. We have the following lemma.
Lemma 2.5.
Let $x\in\mathfrak{g}$.
  1. (1)
    We have
    $J_{G}(x)=G(\mathfrak{z}(\mathfrak{g}^{x_{s}})^{reg}+x_{n}).$
  2. (2)
    $J_{G}(x)$ is irreducible in $\mathfrak{g}$.
  3. (3)
    $J_{G}(x)$ is locally closed in $\mathfrak{g}$, so it is a subvariety of $\mathfrak{g}$.
To a Jordan class $J$, let $x\in J$, and $\mathfrak{l}=\mathfrak{g}^{x_{s}}$. Then $\mathfrak{l}$ is a Levi subalgebra of $\mathfrak{g}$. Let ${\mathbb{O}}_{\mathfrak{l}}$ be the nilpotent orbit in $\mathfrak{l}$ of $x_{n}$. The pair $(\mathfrak{l},{\mathbb{O}}_{\mathfrak{l}})$ does not depend on $x\in J$ up to $G$-conjugacy, and there is a one-to-one correspondence between the set of pairs $(\mathfrak{l},{\mathbb{O}}_{\mathfrak{l}})$ and the set of Jordan classes.
Let $n\in{\mathbb{N}}$. Denote
$\ {\mathfrak{g}}^{(n)}=\{x\in\mathfrak{g}|\dim{\mathfrak{g}}^{x}=n\}.$
Definition 2.6.
For $n\in{\mathbb{N}}$, an irreducible component of $\mathfrak{g}^{(n)}$ is called a sheet of $\mathfrak{g}$.
A sheet of $\mathfrak{g}$ is a disjoint union of Jordan classes. So a sheet ${\mathbb{S}}$ contains a unique dense open Jordan class $J$. The datum $(\mathfrak{l},{\mathbb{O}}_{\mathfrak{l}})$ of the Jordan class $J$ is called the datum of ${\mathbb{S}}$. Then
$\overline{\mathbb{S}}=\overline{J},\qquad{\mathbb{S}}=(\overline{J})^{reg}.$
Let ${\mathbb{S}}$ be a sheet with datum $(\mathfrak{l},{\mathbb{O}}_{\mathfrak{l}})$, then the induced nilpotent orbit ${\rm Ind}_{\mathfrak{l}}^{\mathfrak{g}}({\mathbb{O}}_{\mathfrak{l}})$ of $\mathfrak{g}$ from ${\mathbb{O}}_{\mathfrak{l}}$ in ${\mathfrak{l}}$ is the unique orbit contained in ${\mathbb{S}}$. The rank of ${\mathbb{S}}$ with datum $(\mathfrak{l},{\mathbb{O}}_{\mathfrak{l}})$ is by definition
$r({\mathbb{S}}):=\dim{\mathbb{S}}-\dim({\rm Ind}_{\mathfrak{l}}^{\mathfrak{g}}({\mathbb{O}}_{\mathfrak{l}}))=\dim\mathfrak{z}({\mathfrak{l}}).$ (2.4)
A sheet ${\mathbb{S}}$ with datum $(\mathfrak{l},{\mathbb{O}}_{\mathfrak{l}})$ is called a Dixmier sheet if ${\mathbb{O}}_{\mathfrak{l}}=\{0\}$.
2.7 Affine $W$-algebras
Let $f$ be a nilpotent element of $\mathfrak{g}$. By the Jacobson-Morozov Theorem, there is an ${sl}_{2}$-triple $(e,h,f)$ of $\mathfrak{g}$. Recall that the Slodowy slice $\mathscr{S}_{f}$ is the affine space $f+\mathfrak{g}^{e}$. It has a natural Poisson structure induced from that of $\mathfrak{g}^{*}$ [GG02].
The embedding ${\rm span}_{\mathbb{C}}\{e,h,f\}\cong{sl}_{2}\hookrightarrow\mathfrak{g}$ exponentiates to a homomorphism $SL_{2}\to G$. By restriction to the one-dimensional torus consisting of diagonal matrices, we obtain a one-parameter subgroup $\rho\colon{\mathbb{C}}^{*}\to G$. For $t\in{\mathbb{C}}^{*}$ and $x\in\mathfrak{g}$, set
$\tilde{\rho}(t)x:=t^{2}\rho(t)(x).$
We have $\tilde{\rho}(t)f=f$, and the ${\mathbb{C}}^{*}$-action of $\tilde{\rho}$ stabilizes $\mathscr{S}_{f}$. Moreover, it contracts to $f$ on $\mathscr{S}_{f}$, that is, for all $x\in\mathfrak{g}^{e}$,
$\lim_{t\to 0}\tilde{\rho}(t)(f+x)=f.$
The following proposition is well-known.
Proposition 2.7 ([Slo80, Pre02, CM16]).
The morphism
$\theta_{f}\colon G\times\mathscr{S}_{f}\longrightarrow\mathfrak{g},\quad(g,x)\longmapsto g.x$
is smooth onto a dense open subset of $\mathfrak{g}^{*}$.
Let $W^{k}(\mathfrak{g},f)$ be the affine $W$-algebra associated with a nilpotent element $f$ of $\mathfrak{g}$ defined by the generalized quantized Drinfeld-Sokolov reduction:
$W^{k}(\mathfrak{g},f)=H^{0}_{DS,f}(V^{k}(\mathfrak{g})).$
Here $H^{\bullet}_{DS,f}(M)$ denotes the BRST cohomology of the generalized quantized Drinfeld-Sokolov reduction associated with $f\in\mathcal{N}(\mathfrak{g})$ with coefficients in a $V^{k}(\mathfrak{g})$-module $M$ [KRW03]. Recall that we have a natural isomorphism $R_{W^{k}(\mathfrak{g},f)}\cong{\mathbb{C}}[\mathscr{S}_{f}]$ of Poisson algebras [DSK06, Ara15a], so that
$X_{W^{k}(\mathfrak{g},f)}=\mathscr{S}_{f}.$
We write $W_{k}(\mathfrak{g},f)$ for the unique simple (graded) quotient of $W^{k}(\mathfrak{g},f)$. Then $X_{W_{k}(\mathfrak{g},f)}$ is a ${\mathbb{C}}^{*}$-invariant Poisson subvariety of the Slodowy slice $\mathscr{S}_{f}$.
Let $\mathcal{O}_{k}$ be the category $\mathcal{O}$ of $\widehat{\mathfrak{g}}$ at level $k$. We have a functor
$\mathcal{O}_{k}\longrightarrow W^{k}(\mathfrak{g},f)\operatorname{-Mod},\quad M\longmapsto H^{0}_{DS,f}(M),$
where $W^{k}(\mathfrak{g},f)\operatorname{-Mod}$ denotes the category of $W^{k}(\mathfrak{g},f)$-modules.
The full subcategory of $\mathcal{O}_{k}$ consisting of objects $M$ on which $\mathfrak{g}$ acts locally finitely is denoted by $\operatorname{KL}_{k}$. Note that both $V^{k}(\mathfrak{g})$ and $L_{k}(\mathfrak{g})$ are objects of $\operatorname{KL}_{k}$.
Theorem 2.8 ([Ara16]).
Let $f_{\theta}$ be a root vector of the highest root $\theta$ of $\mathfrak{g}$. Then
  1. (1)
    $H_{DS,f_{\theta}}^{i}(M)=0$ for all $i\neq 0$, $M\in{\mathcal{O}}_{k}$. In particular, the functor
    $\mathcal{O}_{k}\longrightarrow W^{k}(\mathfrak{g},f_{\theta})\operatorname{-Mod},\quad M\longmapsto H^{0}_{DS,f_{\theta}}(M),$
    is exact.
  2. (2)
    $H^{0}_{DS,f_{\theta}}(L(\lambda))\neq 0$ if and only if $\lambda(\alpha_{0}^{\vee})\notin{\mathbb{Z}}_{\geqslant 0}$, where $\alpha_{0}^{\vee}=-\theta^{\vee}+K$. If this is the case, $H^{0}_{DS,f_{\theta}}(L(\lambda))$ is a simple $W^{k}(\mathfrak{g},f_{\theta})$-module.
Theorem 2.9 ([Ara15a]).
Let $k,f$ be arbitrary. Then
  1. (1)
    $H_{DS,f}^{i}(M)=0$ for all $i\neq 0$, $M\in\operatorname{KL}_{k}$. In particular, the functor
    $\operatorname{KL}_{k}\longrightarrow W^{k}(\mathfrak{g},f)\operatorname{-Mod},\quad M\longmapsto H^{0}_{DS,f}(M),$
    is exact.
  2. (2)
    For any quotient $V$ of $V^{k}(\mathfrak{g})$,
    $X_{H^{0}_{DS,f}(V)}=X_{V}\cap\mathscr{S}_{f}.$
    In particular, $H_{DS,f}^{0}(V)\neq 0$ if and only if $\overline{G.f}\subset X_{V}$.
3 Maximal ideals of $V^{k}(sl_{3})$ for $k+n=\frac{n-1}{q}$, $n\geqslant 3$, $q\geqslant 2$, $(q,n-1)=1$
Let $\mathfrak{g}$, $\widehat{\mathfrak{g}}$ and $V^{k}(\mathfrak{g})$ be the same as in Section 2. We first recall the following result from [GK07].
Theorem 3.1 ([GK07]).
For $k\in{\mathbb{C}}$, $V^{k}(\mathfrak{g})$ is not simple if and only if
$r(k+h^{\vee})\in{\mathbb{Q}}_{\geqslant 0}\backslash\left\{\frac{1}{m}\mid m\in{\mathbb{Z}}_{\geqslant 1}\right\},$ (3.1)
where $r$ is the lacing number of $\mathfrak{g}$.
By Theorem 3.1, $V^{k}(sl_{n})$ is not simple if and only if $k=-n+\frac{p}{q}$, where $p\in{\mathbb{Z}}_{\geqslant 2}$, $q\in{\mathbb{Z}}_{\geqslant 1}$, and $(p,q)=1$.
In this section, we will prove that, for $k=-3+\frac{2}{2m+1}$, $m\in{\mathbb{Z}}_{\geqslant 1}$, the maximal ideal of $V^{k}(sl_{3})$ is generated by two singular vectors of conformal weight $3(2m+1)$, and for $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$, the maximal ideal of $V^{k}(sl_{n})$ is generated by a singular vector of conformal weight $2q$. We reach our results by using the Kashiwara-Tanisaki characters [KT00] (see also [Fie06]), the character formulas for the $k=-1$ case given in [KW18, BKK24], and the results established in [AP08, AKMF+20, AM17].
3.1 Kashiwara-Tanisaki Characters
In this subsection, we will introduce the Kashiwara-Tanisaki characters following [KT00]. Let $\Delta$, $\widehat{\Delta}$ be the root systems of $\mathfrak{g}$ and $\widehat{\mathfrak{g}}$ with respect to $\mathfrak{h}$ and $\widehat{\mathfrak{h}}$, respectively. Denote by $\delta$ the positive imaginary root such that any imaginary root is an integral multiple of $\delta$. Let $W$ and $\widehat{W}$ be the Weyl groups of $\mathfrak{g}$ and $\widehat{\mathfrak{g}}$, and $Q$ and $\widehat{Q}$ the root lattices of $\mathfrak{g}$ and $\widehat{\mathfrak{g}}$, respectively. Let $\{\alpha_{1},\dots,\alpha_{l}\}$ be the simple root system of $\mathfrak{g}$, and $\{\alpha_{0},\alpha_{1},\dots,\alpha_{l}\}$ the simple root system of $\widehat{\mathfrak{g}}$. For a real root $\alpha\in\widehat{\Delta}$, set $\alpha^{\vee}=2\alpha/(\alpha|\alpha)$. Let $\rho\in{\mathfrak{h}}^{*}$ and $\widehat{\rho}\in\widehat{\mathfrak{h}}^{*}$ be such that
[…]
Recall that the twisted (or shifted) action of $\widehat{W}$ on $\widehat{\mathfrak{h}}^{*}$ is defined as follows
$w\circ\lambda=w(\lambda+\widehat{\rho})-\widehat{\rho},$
where $w\in\widehat{W}$ and $\lambda\in\widehat{\mathfrak{h}}^{*}$. Notice that if $w\in W$, $w\circ\lambda=w(\lambda+\rho)-\rho$.
For $\lambda\in\widehat{\mathfrak{h}}^{*}$, set
$\widehat{\Delta}(\lambda)=\{\alpha\in\widehat{\Delta}^{re}|(\lambda+\widehat{\rho}|\alpha^{\vee})\in{\mathbb{Z}}\},$
$\widehat{\Delta}_{0}(\lambda)=\{\alpha\in\widehat{\Delta}^{re}|(\lambda+\widehat{\rho}|\alpha^{\vee})=0\}.$
Notice that $\widehat{\Delta}(\lambda)$ and $\widehat{\Delta}_{0}(\lambda)$ are subsystems of $\widehat{\Delta}^{re}$. Denote the set of positive roots, the set of negative roots, the set of simple roots and the Weyl group for $\widehat{\Delta}(\lambda)$ by $\widehat{\Delta}^{+}(\lambda)$, $\widehat{\Delta}^{-}(\lambda)$, $\widehat{\Pi}(\lambda)$ and $\widehat{W}(\lambda)$, respectively. Denote those for $\widehat{\Delta}_{0}(\lambda)$ by $\widehat{\Delta}^{+}_{0}(\lambda)$, $\widehat{\Delta}^{-}_{0}(\lambda)$, $\widehat{\Pi}_{0}(\lambda)$ and $\widehat{W}_{0}(\lambda)$.
For a real root $\alpha\in\widehat{\Delta}$, denote by $s_{\alpha}\in\widehat{W}$ the corresponding reflection. Then $\widehat{\Pi}(\lambda)$ is the set of $\alpha\in\widehat{\Delta}^{+}(\lambda)$ such that $s_{\alpha}(\widehat{\Delta}^{+}(\lambda)\backslash\{\alpha\})=\widehat{\Delta}^{+}(\lambda)\backslash\{\alpha\}$, and $(\widehat{W}(\lambda),S(\lambda))$ is a Coxeter group, where $S(\lambda)=\{s_{\alpha}:\alpha\in\widehat{\Pi}(\lambda)\}$ [KW89], [KT00].
For $w\in\widehat{W}(\lambda)$, denote by $l_{\lambda}(w)$ the length of $w$. Denote the Bruhat ordering of $\widehat{W}(\lambda)$ by $\geqslant_{\lambda}$. For $y,w\in\widehat{W}(\lambda)$, denote by $P^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the corresponding Kazhdan-Lusztig polynomial [KL79], and by $Q^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the inverse Kazhdan-Lusztig polynomial defined by
$\sum\limits_{x\leqslant_{\lambda}y\leqslant_{\lambda}z}(-1)^{l_{\lambda}(y)-l_{\lambda}(x)}Q^{\lambda}_{x,y}(q)P^{\lambda}_{y,z}(q)=\delta_{x,z},$
for any $x,z\in\widehat{W}(\lambda)$. Set
${\mathcal{C}}=\{\lambda\in\widehat{\mathfrak{h}}^{*}|(\delta|\lambda+\widehat{\rho})\neq 0\},$
${\mathcal{C}}^{+}=\{\lambda\in{\mathcal{C}}|(\lambda+\widehat{\rho}|\alpha^{\vee})\geqslant 0,\ {\rm for\ any}\ \alpha\in\widehat{\Delta}^{+}(\lambda)\},$
and
${\mathcal{C}}^{-}=\{\lambda\in{\mathcal{C}}|(\lambda+\widehat{\rho}|\alpha^{\vee})\leqslant 0,\ {\rm for\ any}\ \alpha\in\widehat{\Delta}^{+}(\lambda)\}.$
Let $\lambda\in{\mathcal{C}}$. Then $\widehat{W}_{0}(\lambda)$ is a finite group [KT00].
For $\lambda\in\widehat{\mathfrak{h}}^{*}$, let $M(\lambda)$ (resp. $L(\lambda)$) be the Verma module (resp. simple module) of $\widehat{\mathfrak{g}}$ with highest weight $\lambda$. We have the following result from [KT00].
Theorem 3.2.
[KT00]
  1. (1)
    Let $\lambda\in{\mathcal{C}}^{+}$, then for any $w\in\widehat{W}(\lambda)$ which is the longest element of $w\widehat{W}_{0}(\lambda)$,
    ${\rm ch}(L(w\circ\lambda))=\sum\limits_{w\leqslant_{\lambda}y\in\widehat{W}(\lambda)}(-1)^{l_{\lambda}(y)-l_{\lambda}(w)}Q^{\lambda}_{w,y}(1){\rm ch}(M(y\circ\lambda)).$
  2. (2)
    Let $\lambda\in{\mathcal{C}}^{-}$, then for any $w\in\widehat{W}(\lambda)$ which is the shortest element of $w\widehat{W}_{0}(\lambda)$,
    ${\rm ch}(L(w\circ\lambda))=\sum\limits_{w\geqslant_{\lambda}y\in\widehat{W}(\lambda)}(-1)^{l_{\lambda}(w)-l_{\lambda}(y)}P^{\lambda}_{y,w}(1){\rm ch}(M(y\circ\lambda)).$
[…]
In this subsection, we assume that $\mathfrak{g}=sl_{n}({\mathbb{C}})$, and $k=-n+\frac{n-1}{q}$, $q\geqslant 1$, $(q,n-1)=1$. It is obvious that
$\widehat{\Pi}(k\Lambda_{0})=\{\beta_{0}=q\delta-\theta,\alpha_{1},\cdots,\alpha_{n-1}\}.$
Since $(k\Lambda_{0}+\widehat{\rho}|\beta_{0})=0$, it follows that
$\widehat{\Pi}_{0}(k\Lambda_{0})=\{\beta_{0}\},$
and
$\widehat{W}_{0}(k\Lambda_{0})=\{e,s_{\beta_{0}}\}.$
In particular, if $q=1$, then $k=-1$ and $\beta_{0}=\alpha_{0}$. It follows that
$\widehat{W}(-\Lambda_{0})=\widehat{W},\ \widehat{W}_{0}(-\Lambda_{0})=\{e,s_{\alpha_{0}}\}.$
Let $S=\{s_{\alpha_{0}},s_{\alpha_{1}},\cdots,s_{\alpha_{n-1}}\}$, then $(\widehat{W},S)$ is a Coxeter group. We have the following lemma.
Lemma 3.3.
For any $k=-n+\frac{n-1}{q}$, $q\geqslant 1$, $n\geqslant 3$, we have
$(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))\cong(\widehat{W},S)$
as Coxeter groups by $\sigma:(\widehat{W},S)\to(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$ such that
$\sigma s_{\alpha_{0}}=s_{\beta_{0}},\sigma(s_{\alpha_{i}})=s_{\alpha_{i}},\ i=1,2,\cdots,n-1.$
Proof.
Recall that the defining relations of $(\widehat{W},S)$ are [Kac90]
$s_{\alpha_{i}}^{2}=1,\ i=0,1,\cdots,n-1,\ (s_{\alpha_{i}}s_{\alpha_{j}})^{m_{ij}}=1\ i,j=0,1,\cdots,n-1,$
such that
$m_{01}=m_{10}=m_{0,n-1}=m_{n-1,0}=3,\ m_{0j}=m_{j0}=2,\ j=2,\cdots,n-2,$
[…]
Notice that $S=\{s_{\alpha_{0}},s_{\alpha_{1}},\cdots,s_{\alpha_{n-1}}\}$, $S(k\Lambda_{0})=\{s_{\beta_{0}},s_{\alpha_{1}},\cdots,s_{\alpha_{n-1}}\}$. Thus it is enough to prove that
$(s_{\beta_{0}}s_{\alpha_{1}})^{3}=(s_{\alpha_{1}}s_{\beta_{0}})^{3}=(s_{\beta_{0}}s_{\alpha_{n-1}})^{3}=(s_{\alpha_{n-1}}s_{\beta_{0}})^{3}=1,$ (3.2)
and
$(s_{\beta_{0}}s_{\alpha_{j}})^{2}=(s_{\alpha_{j}}s_{\beta_{0}})^{2}=1,\ j=2,\cdots,n-2.$ (3.3)
We only check $(s_{\beta_{0}}s_{\alpha_{1}})^{3}=1$. The computation for the other cases in (3.2)-(3.3) is similar. Notice that
$s_{\beta_{0}}s_{\alpha_{1}}s_{\beta_{0}}=s_{s_{\beta_{0}}(\alpha_{1})}=s_{q\delta-\theta+\alpha_{1}},$
and
$s_{\alpha_{1}}s_{q\delta-\theta+\alpha_{1}}s_{\alpha_{1}}=s_{s_{\alpha_{1}}(q\delta-\theta+\alpha_{1})}=s_{q\delta-\theta}=s_{\beta_{0}}.$
Then we have
$(s_{\beta_{0}}s_{\alpha_{1}})^{3}=s_{\beta_{0}}s_{\alpha_{1}}s_{\beta_{0}}s_{\alpha_{1}}s_{\beta_{0}}s_{\alpha_{1}}=s_{\beta_{0}}s_{\alpha_{1}}s_{q\delta-\theta+\alpha_{1}}s_{\alpha_{1}}=s_{\beta_{0}}s_{\beta_{0}}=1.$
For $\gamma\in\mathfrak{h}^{*}$, let $t_{\gamma}\in{\rm End}\widehat{\mathfrak{h}}^{*}$ be defined by [Kac90]
$t_{\gamma}(\lambda)=\lambda+(\lambda|\delta)\gamma-((\lambda|\gamma)+\frac{1}{2}(\lambda|\delta)(\gamma|\gamma))\delta,$
for $\lambda\in\widehat{\mathfrak{h}}^{*}$. Recall from [Kac90], for any $x\in\widehat{W}$, there exists $y\in W$ and $\gamma\in Q$ such that $x=yt_{\gamma}$. We have the following lemma.
Lemma 3.4.
For $\gamma\in Q$, $\sigma(t_{\gamma})=t_{q\gamma}$.
Proof.
We first consider $\gamma=\theta$. Notice that for $\lambda\in\widehat{\mathfrak{h}}^{*}$,
$s_{\beta_{0}}s_{\theta}(\lambda)=s_{\beta_{0}}(\lambda-(\lambda|\theta)\theta)=\lambda+(\lambda|\delta)q\theta-((\lambda|\theta)q+((\lambda|\delta)q^{2})\delta=t_{q\theta}(\lambda).$
This deduces that
$s_{\beta_{0}}s_{\theta}=t_{q\theta}.$
In particular, if $q=1$, $s_{\alpha_{0}}s_{\theta}=t_{\theta}$ (see also [Kac90]). So by Lemma 3.3,
$\sigma(t_{\theta})=t_{q\theta}.$ (3.4)
Notice that
$s_{\alpha_{1}}t_{\theta}s_{\alpha_{1}}=t_{s_{\alpha_{1}}(\theta)}=t_{\theta-\alpha_{1}}$
and
$s_{\alpha_{1}}t_{q\theta}s_{\alpha_{1}}=t_{s_{\alpha_{1}}(q\theta)}=t_{q(\theta-\alpha_{1})}.$
Thus
$\sigma(t_{\theta-\alpha_{1}})=t_{q(\theta-\alpha_{1})}.$ (3.5)
Considering $s_{\alpha_{2}}t_{\theta-\alpha_{1}}s_{\alpha_{2}}$ and $s_{\alpha_{2}}t_{q(\theta-\alpha_{1})}s_{\alpha_{2}}$, we deduce that
$\sigma(t_{\theta-\alpha_{1}-\alpha_{2}})=t_{q(\theta-\alpha_{1}-\alpha_{2})}.$ (3.6)
Similarly, we have for $14\leqslant j\leqslant n-$ 1,
$\sigma(t_{\alpha_{j}+\cdots\alpha_{n-1}})=t_{q(\alpha_{j}+\cdots\alpha_{n-1})}.$ (3.7)
Notice that for $\alpha,\beta\in Q$, $t_{\alpha}t_{\beta}=t_{\alpha+\beta}$. Then the lemma follows from (3.4)-(3.7). ∎
Let $\Lambda_{i}\in\widehat{\mathfrak{h}}^{*}$ be such that
$\Lambda_{i}(\alpha_{j}^{\vee})=\delta_{ij},\ i,j=0,1,\cdots,n-1,$
We have the following character formulas from [KW18], [BKK24].
Theorem 3.5 ([KW18, BKK24]).
Let $n\geqslant 3$, and let $L(\Lambda)$ be an irreducible level $k=-1$ $\widehat{sl}_{n}$-module with highest weight
$\Lambda=-(1+s)\Lambda_{0}+s\Lambda_{n-1}\ ({\rm resp.}\ \Lambda=-(1+s)\Lambda_{0}+s\Lambda_{1}),\ s\in{\mathbb{Z}}_{\geqslant 0}.$
Then the character of $L(\Lambda)$ is given by the following formula:
$\widehat{R}{\rm ch}L(\Lambda)=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1}(resp.\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}({\Lambda+\widehat{\rho}})-\widehat{\rho}}.$
where
$\widehat{R}=\prod_{\alpha\in\widehat{\Delta}_{+}}(1-e^{-\alpha})^{{\rm dim}\widehat{\mathfrak{g}}_{\alpha}}.$
In particular, if $\Lambda=-\Lambda_{0}$,
$\widehat{R}{\rm ch}L(-\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
We also need the following results from [KK79], [DGK82], [Kum87], [KW89].
Lemma 3.6.
Let $\lambda\in{\mathcal{C}}^{+}$, and $w\in\widehat{W}(\lambda)$. If $L(\mu)$ is a subquotient of $M(w\circ\lambda)$, then there exists $y\in\widehat{W}(\lambda)$ such that $\mu=yw\circ\lambda$.
Recall the following results from [AP08], [AKMF+20], and [AM17].
Theorem 3.7.
  1. (1)
    [AP08] Let
    $\begin{array}[]{ll}u^{1}=&[-e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)+e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{2}(-1)+e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{2}}(-1)]{\bf 1},\\ \\ u^{2}=&[e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}+e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)\\ &-2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-2)-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}}(-1)]{\bf 1}.\end{array}$
    Then $u^{1},u^{2}$ are singular vectors of $V^{-1}(sl_{3})$.
  2. (2)
    [AP08] Let
    $u=e_{\theta}(-1)e_{\alpha_{2}+\cdots\alpha_{n-2}}(-1){\bf 1}-e_{\alpha_{1}+\cdots\alpha_{n-2}}(-1)e_{\alpha_{2}+\cdots\alpha_{n-1}}(-1){\bf 1}.$
    Then $u$ is a singular vector of $V^{-1}(sl_{n})$, $n\geqslant 4$.
  3. (3)
    [AKMF+20] The ideal of $V^{-1}(sl_{3})$ generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$.
  4. (4)
    [AM17] The ideal generated by $u$ in (2) is the maximal ideal of $V^{-1}(sl_{n})$, $n\geqslant 4$.
We are now in a position to state the main result of this section.
Theorem 3.8.
Let $k=-n+\frac{n-1}{q}$, $n\geqslant 3$, $q\geqslant 2$, $(n-1,q)=1$. Then
$\widehat{R}{\rm ch}L_{k}(sl_{n})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{q\gamma}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$ (3.8)
Furthermore, we have
  1. (1)
    If $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, then the maximal ideal of $V^{k}(sl_{3})$ is generated by two singular vectors $v^{1}$ and $v^{2}$ of weights $k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2}$ and $k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2}$, respectively.
  2. (2)
    If $k=-n+\frac{n-1}{q}$ satisfying $n\geqslant 4$, $q\geqslant 2$, and $(q,n-1)=1$, then the maximal ideal of $V^{k}(sl_{n})$ is generated by a singular vector $w^{1}$ of weight $k\Lambda_{0}-2q\delta+\alpha_{1}+2\alpha_{2}+\cdots+2\alpha_{n-2}+\alpha_{n-1}$.
Proof.
Let $\mathfrak{g}=sl_{n}$, $n\geqslant 3$. Notice that if $q=1$, then $k=-1$ and $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))=(\widehat{W},S)$. So $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))\cong(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2,(q,n-1)=1$. By Lemma 3.3, and the definitions of $P^{\lambda}_{y,w}(q)$ and $Q^{\lambda}_{y,w}(q)$ [KL79], we have for $y,w\in\widehat{W}$,
$w\geqslant_{-\Lambda_{0}}y\ {\rm if\ and\ if}\ \sigma(w)\geqslant_{k\Lambda_{0}}\sigma(y),$ (3.9)
$l_{-\Lambda_{0}}(y)=l_{k\Lambda_{0}}\sigma(y),\ l_{-\Lambda_{0}}(w)=l_{k\Lambda_{0}}\sigma(w),$ (3.10)
and
$P^{-\Lambda_{0}}_{y,w}(1)=P^{k\Lambda_{0}}_{\sigma(y),\sigma(w)}(1),\ Q^{-\Lambda_{0}}_{y,w}(1)=Q^{k\Lambda_{0}}_{\sigma(y),\sigma(w)}(1).$ (3.11)
It is obvious that $-\Lambda_{0},k\Lambda_{0}\in{\mathcal{C}}^{+}$, for $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$. Notice that for any $w\in\widehat{W}(-\Lambda_{0})$ which is the longest element of $w\widehat{W}_{0}(-\Lambda_{0})$, $\sigma(w)$ is the longest element of $\sigma(w)\widehat{W}_{0}(k\Lambda_{0})$. Then by Theorem 3.2 and (3.9)-(3.11) ( see also Theorem 11 of [Fie06] from the category point of view),
${\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1){\rm ch}(M(\sigma(y)\circ(k\Lambda_{0}))),$ (3.12)
which is equivalent to
$\widehat{R}{\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1)e^{\sigma(y)\circ k\Lambda_{0}}.$ (3.13)
Then by Lemma 3.3, Lemma 3.4, and Theorem 3.5,
$\widehat{R}{\rm ch}L(k\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{q\gamma}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
Furthermore, (3.12) together with Lemma 3.6 means that $L(y\circ(-\Lambda_{0}))$ for some $y\in\widehat{W}(-\Lambda_{0})$ is isomorphic to a subquotient of $M(w\circ(-\Lambda_{0}))$ if and only if $L(\sigma(y)\circ k\Lambda_{0})$ is a subquotient of $M(\sigma(w)\circ k\Lambda_{0})$.
Case 1   $n=3$, $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$. Let $u^{1}$ and $u^{2}$ be the singular vectors given in (1) of Theorem 3.7. The weights of $u^{1}$ and $u^{2}$ with respect to $\widehat{\mathfrak{h}}$ are $-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2}$ and $-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}$, respectively. It is easily checked that
$s_{\alpha_{2}}\circ(t_{\alpha_{1}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2},$ (3.14)
and
$s_{\alpha_{1}}\circ(t_{\alpha_{2}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}.$ (3.15)
Thus $V^{k}(sl_{3})$ has two singular vectors $v^{1}$ and $v^{2}$ of wights $s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$ and $s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$, respectively. It is easy to check that
$s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2},$ (3.16)
and
$s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2},$ (3.17)
By (3) of Theorem 3.7, the ideal generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$. It follows that the ideal generated by $v^{1}$ and $v^{2}$ is the maximal ideal of $V^{k}(sl_{3})$. We complete the proof of (1).
Case 2  $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$. Let $u$ be the singular vector in (2) of Theorem 3.7. The weight of $u$ with respect to the Cartan subalgebra $\widehat{\mathfrak{h}}$ is $-\Lambda_{0}-2\delta+\theta+\beta$, where $\beta=\alpha_{2}+\cdots+\alpha_{n-2}$. It is easily checked that
$-\Lambda_{0}-2\delta+\theta+\beta=(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}})\circ(t_{-\beta}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}).$
Then $V^{k}(sl_{n})$ has a singular vector $w^{1}$ of weight $(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}})\circ(t_{-q\beta}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$. It can be checked directly that
$(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}})\circ(t_{-q\beta}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-2q\delta+\theta+\beta.$
By (4) of Theorem 3.7, the ideal generated by $u$ is the maximal ideal of $V^{-1}(sl_{n})$. It follows that the ideal generated by $w^{1}$ is the maximal ideal of $V^{k}(sl_{n})$. ∎
Remark 3.9.
The method here can be used to the subregular case for $\mathfrak{g}$ of other types.
4 Varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$
Let $\mathfrak{g}$ and $V^{k}(\mathfrak{g})$ be as above in Section 3. For $x\in V^{k}(\mathfrak{g})$, we denote by $\bar{x}$ the image of $x$ in $R_{V^{k}(\mathfrak{g})}$. Recall that $R_{V^{k}(\mathfrak{g})}\cong{\mathbb{C}}[\mathfrak{g}^{*}]$ by
$\overline{x_{1}(-1)\cdots x_{n}(-1){\bf 1}}\mapsto x_{1}\cdots x_{n},$
for $x_{1},\cdots,x_{n}\in\mathfrak{g}$. As in Section 2, we denote by ${\mathcal{I}}_{k}$ the maximal ideal of $V^{k}(\mathfrak{g})$, then
$L_{k}(\mathfrak{g})=V^{k}(\mathfrak{g})/{\mathcal{I}}_{k}$
and
$R_{L_{k}(\mathfrak{g})}={\mathbb{C}}[\mathfrak{g}^{*}]/\overline{{\mathcal{I}}_{k}},$
where $\overline{{\mathcal{I}}_{k}}$ is the image of $\mathcal{I}_{k}$ in $R_{V^{k}(\mathfrak{g})}$.
4.1 The associated variety of $L_{-1}(sl_{3})$
Let $\Delta$ be the root system of $sl_{3}({\mathbb{C}})$ and
$\{h_{1},h_{2},e_{\alpha_{1}},e_{\alpha_{2}},e_{\alpha_{1}+\alpha_{2}},f_{\alpha_{1}},f_{\alpha_{2}},f_{\alpha_{1}+\alpha_{2}}\}$
be a Chevalley basis of $sl_{3}({\mathbb{C}})$ with structure constants $c_{\alpha,\beta}$, for $\alpha,\beta\in\Delta$ such that
$c_{\alpha_{1},\alpha_{2}}=1.$
Then we have
$c_{\alpha_{2},-\alpha_{1}-\alpha_{2}}=c_{-\alpha_{1}-\alpha_{2},\alpha_{1}}=1.$
Recall from Theorem 3.7 [AP08] that
$\begin{array}[]{ll}u^{1}=&[-e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)+e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{2}(-1)+e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{2}}(-1)]{\bf 1},\\ \\ u^{2}=&[e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}+e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)\\ &-2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-2)-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}}(-1)]{\bf 1}.\end{array}$
are singular vectors of $V^{-1}(sl_{3})$. Then we have
Lemma 4.1.
There exist $x,y\in{\mathcal{I}}_{-1}$ such that
$\begin{array}[]{ll}x\equiv&-e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)h_{2}(-1){\bf 1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1){\bf 1}\\ \\ &-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}+\alpha_{2}}(-1){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)(2h_{1}+h_{2})(-1){\bf 1}\\ \\ &+2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{2}}(-1){\bf 1}({\rm mod}~{}F^{1}(V^{-1}(sl_{3})))\\ \\ y\equiv&h_{1}(-1)h_{2}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{3}}(-1){\bf 1}\\ \\ &-3e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1)f_{\alpha_{2}}(-1){\bf 1}-e_{\alpha_{1}}(-1)f_{\alpha_{1}}(-1)h_{1}(-1){\bf 1}\\ \\ &+e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{2}}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}\\ \\ &-e_{\alpha_{2}}(-1)f_{\alpha_{2}}(-1)h_{2}(-1){\bf 1}\ ({\rm mod}~{}F^{1}(V^{-1}(sl_{3}))).\end{array}$
Proof.
Let $x=f_{\alpha_{1}}(0)u^{1}$, and $y=f_{\alpha_{1}}(0)f_{\alpha_{1}+\alpha_{2}}(0)u^{1}\in{\mathcal{I}}_{-1}$. Then the lemma follows. ∎
Let
$\begin{array}[]{ll}p_{1}=&-e_{\alpha_{1}}^{2}e_{\alpha_{2}}+e_{\alpha_{1}}e_{\alpha_{1}+\alpha_{2}}h_{2}+e_{\alpha_{1}+\alpha_{2}}^{2}f_{\alpha_{2}},\\ \\ p_{2}=&e_{\alpha_{1}}e_{\alpha_{2}}^{2}+e_{\alpha_{2}}e_{\alpha_{1}+\alpha_{2}}h_{1}-e_{\alpha_{1}+\alpha_{2}}^{2}f_{\alpha_{1}},\\ \\ p_{3}=&-e_{\alpha_{1}+\alpha_{2}}h_{1}h_{2}-e_{\alpha_{1}}e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{1}}-e_{\alpha_{1}+\alpha_{2}}^{2}f_{\alpha_{1}+\alpha_{2}}+e_{\alpha_{1}}e_{\alpha_{2}}(2h_{1}+h_{2})+2e_{\alpha_{2}}e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{2}},\\ \\ p_{4}=&h_{1}h_{2}(h_{1}+h_{2})+e_{\alpha_{1}}e_{\alpha_{2}}f_{\alpha_{1}+\alpha_{3}}-3e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{1}}f_{\alpha_{2}}-e_{\alpha_{1}}f_{\alpha_{1}}h_{1},\\ \\ &+e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{1}+\alpha_{2}}(h_{1}+h_{2})-e_{\alpha_{2}}f_{\alpha_{2}}h_{2}.\end{array}$
Then by Lemma 4.1, $p_{i}\in\overline{{\mathcal{I}}_{-1}}$, $1\leqslant i\leqslant 4$.
Theorem 4.2.
Denote $\lambda=h_{1}-h_{2}$. Then
$X_{L_{-1}(sl_{3})}=\overline{{\mathbb{S}}_{\mathfrak{l}_{1}}}=\overline{G.{\mathbb{C}}^{*}\lambda}=G.{\mathbb{C}}^{*}\lambda\cup\overline{{\mathcal{O}}_{\min}},$
which is the closure of the Dixmier sheet ${\mathbb{S}}_{\mathfrak{l}_{1}}=G.{\mathbb{C}}^{*}\lambda$, where ${\mathfrak{l}}_{1}=\mathfrak{h}\oplus{\mathbb{C}}e_{\theta}\oplus{\mathbb{C}}f_{\theta}$. In particular, $X_{L_{-1}(sl_{3})}$ is irreducible and
$\dim X_{L_{-1}(sl_{3})}=5,$
where ${\mathcal{O}}_{\min}$ is the minimal non-zero nilpotent orbit of $sl_{3}({\mathbb{C}})$.
Proof.
It is known that for any simple Lie algebra $\mathfrak{g}$ over ${{\mathbb{C}}}$, $\mathfrak{g}$ is the finite disjoint union of its Jordan classes [TY05]. In particular, for $\mathfrak{g}=sl_{3}({\mathbb{C}})$, we have
$\mathfrak{g}=G\cdot({\mathbb{C}}^{*}\lambda)\cup G\cdot({\mathbb{C}}^{*}\lambda+f_{\theta})\cup G\cdot({\mathfrak{h}}^{reg})\cup G\cdot f_{\theta}\cup G\cdot(f_{\alpha_{1}}+f_{\alpha_{2}})\cup\{0\},$
where $G$ is the connected adjoint group of $\mathfrak{g}$, and $\theta=\alpha_{1}+\alpha_{2}$. Notice that
$p_{1}(f_{\alpha_{1}}+f_{\alpha_{2}})=-1,\ p_{3}(t\lambda+f_{\theta})=-t^{2}.$
This implies that $G\cdot(f_{\alpha_{1}}+f_{\alpha_{2}})$ and $G\cdot({\mathbb{C}}^{*}\lambda+f_{\theta})$ could not belong in $X_{L_{-1}(sl_{3})}$. Furthermore, for regular semisimple vector $h$, all $(h_{1}|h),\ (h|h_{2}),(h|h_{1}+h_{2})$ are non-zero. So
$p_{4}(h)=(h|h_{1})(h|h_{2})(h|h_{1}+h_{2})\neq 0.$
We see that for any $h\in{\mathfrak{h}}^{reg}$, $h\notin X_{L_{-1}(sl_{3})}$. We deduce that
$X_{L_{-1}(sl_{3})}\subseteq\overline{G.{\mathbb{C}}^{*}\lambda}=G.{\mathbb{C}}^{*}\lambda\cup\overline{{\mathcal{O}}_{\min}}.$
By (3) of Theorem 3.7, the ideal generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$. This means that
$X_{L_{-1}(sl_{3})}=\overline{G.{\mathbb{C}}^{*}\lambda}=G.{\mathbb{C}}^{*}\lambda\cup\overline{{\mathcal{O}}_{\min}}.$
It follows that $X_{L_{-1}(sl_{3})}$ is irreducible and $\dim X_{L_{-1}(sl_{3})}=5$. ∎
4.2 Associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1},m\geqslant 1$
In this subsection we always assume that $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$.
Recall from Theorem 3.8 that $V^{k}(sl_{3})$ has two singular vectors $v^{1}$ and $v^{2}$ with weights $k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2}$ and $k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2}$, respectively. Then for $i=1,2$,
$e_{\alpha_{1}}(0)v^{i}=e_{\alpha_{2}}(0)v^{i}=f_{\alpha_{1}}(1)v^{i}=f_{\alpha_{2}}(1)v^{i}=0.$
Notice that each element in $V^{k}(sl_{3})$ is a linear combination of elements of the following form:
$z=z^{(+)}z^{(-)}z^{(0)}{\bf 1},$ (4.1)
with
$z^{(+)}:=$ $e_{\alpha_{1}}(-1)^{a_{11}}\cdots e_{\alpha_{1}}(-r_{1})^{a_{1r_{1}}}e_{\alpha_{2}}(-1)^{a_{21}}\cdots e_{\alpha_{2}}(-r_{2})^{a_{2r_{2}}}$
$e_{\alpha_{1}+\alpha_{2}}(-1)^{a_{31}}\cdots e_{\alpha_{1}+\alpha_{2}}(-r_{3})^{a_{3r_{3}}},$
$z^{(-)}:=$ $f_{\alpha_{1}}(-1)^{b_{11}}\cdots f_{\alpha_{1}}(-s_{1})^{b_{1s_{1}}}f_{\alpha_{2}}(-1)^{b_{21}}\cdots f_{\alpha_{2}}(-s_{2})^{b_{2s_{2}}}$
$f_{\alpha_{1}+\alpha_{2}}(-1)^{b_{31}}\cdots f_{\alpha_{1}+\alpha_{2}}(-s_{3})^{b_{3s_{3}}},$
$z^{(0)}:=$ $h_{1}(-1)^{c_{11}}\cdots h_{1}(-t_{1})^{c_{1t_{1}}}h_{2}(-1)^{c_{21}}\cdots h_{2}(-t_{2})^{c_{2t_{2}}},$
where $r_{1},r_{2},r_{3},s_{1},s_{2},s_{3},,t_{1},t_{2}$ are positive integers, and $a_{lp},b_{ln},c_{ij}$, for $l=1,2,3$, $p=1,\ldots,r_{l}$, $n=1,\ldots,s_{l}$, $i=1,2$, $j=1,\ldots,t_{i}$, are nonnegative integers such that at least one of them is non-zero. Recall that
$depth(z^{(+)})=\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{r_{i}-1}(j-1)a_{ij},~{}depth(z^{(-)})=\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{s_{i}-1}(j-1)b_{ij},$
$depth(z^{(0)})=\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{t_{i}-1}(j-1)c_{ij};$
$deg{(z^{(+)})}=\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{r_{i}}a_{ij},~{}deg{(z^{(-)})}=\sum\limits_{i=1}^{3}\sum\limits_{j=1}^{s_{i}}b_{ij},~{}deg{(z^{(0)})}=\sum\limits_{i=1}^{2}\sum\limits_{j=1}^{t_{i}}c_{ij};$
$depth(z)=depth(z^{(+)})+depth(z^{(-)})+depth(z^{(0)});$
$deg(z)=deg(z^{(+)})+deg(z^{(-)})+deg(z^{(0)}).$
Let $V^{1}$ be the subspace of $V^{k}(sl_{3})$ linearly spanned by elements $z=z^{(+)}z^{(-)}z^{(0)}{\bf 1}$ of weight $k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2}$ such that $deg(z^{(0)})\leqslant 6m-2$ or $depth(z^{(0)})\geqslant 1$. Then we may assume that
$\begin{array}[]{ll}v^{1}=&\sum\limits_{i=0}^{6m+1}a_{i}e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{i}h_{2}(-1)^{6m+1-i}{\bf 1}\\ \\ &+\sum\limits_{i=0}^{6m}[x_{i}e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)+y_{i}e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{2}}(-1)+b_{i}e_{\alpha_{1}}(-2)e_{\alpha_{1}+\alpha_{2}}(-1)\\ \\ &~{}~{}~{}+c_{i}e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-2)]h_{1}(-1)^{i}h_{2}(-1)^{6m-i}{\bf 1}\\ \\ &+\sum\limits_{i=0}^{6m-1}[d_{i}e_{\alpha_{1}+\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-2)f_{\alpha_{2}}(-1)+z_{i}e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)\\ \\ &f_{\alpha_{2}}(-1)+l_{i}e_{\alpha_{1}}(-1)^{2}e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1)+n_{i}e_{\alpha_{1}}(-2)e_{\alpha_{1}+\alpha_{2}}(-2)\\ \\ &+k_{i}e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-2)+g_{i}e_{\alpha_{1}}(-1)e_{\alpha_{1}}(-2)e_{\alpha_{2}}(-1)+p_{i}e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{2}}(-2)\\ \\ &+q_{i}e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}+\alpha_{2}}(-1)+m_{i}e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-3)\\ \\ &+r_{i}e_{\alpha_{1}}(-3)e_{\alpha_{1}+\alpha_{2}}(-1)]h_{1}(-1)^{i}h_{2}(-1)^{6m-1-i}{\bf 1}+u^{1},\end{array}$
where $u^{1}\in V^{1}$.
The following result comes from Lemma 3.1 of [AJM21].
Lemma 4.3.
Not all $a_{i},x_{j},0\leqslant i\leqslant 6m+1,0\leqslant j\leqslant 6m$ in $v^{1}$ are zero.
We now have the following lemma.
Lemma 4.4.
[…]
$(6m+1)a_{6m+1}-2a_{6m}-2x_{6m}=0,$ (4.5)
[…]
$(-4+\frac{2}{2m+1})a_{6m+1}-b_{6m}=0,\ (-4+\frac{2}{2m+1})a_{0}+2y_{0}=0,$ (4.7)
[…]
$a_{0}=y_{0}=0,$ (4.9)
[…]
Proof.
We consider $e_{\alpha_{1}}(0)v^{1}$, $e_{\alpha_{2}}(0)v^{1}$, $f_{\alpha_{1}}(1)v^{1}$, and $f_{\alpha_{1}+\alpha_{2}}(1)v^{1}$. By the definition of $V^{1}$, there are no monomials $z=z^{(+)}z^{(-)}z^{(0)}{\bf 1}$ in $e_{\alpha_{1}}(0)u^{1}$, $e_{\alpha_{2}}(0)u^{1}$, $f_{\alpha_{1}}(1)u^{1}$, $f_{\alpha_{2}}(1)u^{1}$, and $f_{\alpha_{1}+\alpha_{2}}(1)u^{1}$ such that $deg(z^{(0)})\geqslant 6m$, $depth(z)=0$, and $z^{(-)}=1$. Then it is easy to deduce that the coefficients of $e_{\alpha_{1}}(-1)^{2}e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{i}h_{2}(-1)^{6m-i}{\bf 1}$ in $e_{\alpha_{1}}(0)v^{1}$, $0\leqslant i\leqslant 6m$, are
[…]
Then (4.3) holds.
The coefficients of $e_{\alpha_{1}+\alpha_{2}}(-1)^{2}h_{1}(-1)^{i}h_{2}(-1)^{6m+1-i}{\bf 1}$ in $e_{\alpha_{2}}(0)v^{1}$, $0\leqslant i\leqslant 6m$, are
[…]
and the coefficient of $e_{\alpha_{1}+\alpha_{2}}(-1)^{2}h_{1}(-1)^{6m+1-i}{\bf 1}$ is $a_{6m+1}$. Then (4.2) follows.
The coefficients of $e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{i}h_{2}(-1)^{6m-i}{\bf 1}$, $10\leqslant i\leqslant 6m-$ 1, in $e_{\alpha_{2}}(0)v^{1}$ are
[…]
and the coefficient of $e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{6m}{\bf 1}$ in $e_{\alpha_{2}}(0)v^{1}$ is
$(6m+1)a_{6m+1}-2a_{6m}-2x_{6m}.$
Then (4.4) and (4.5) hold.
The coefficients of $e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{i}h_{2}(-1)^{6m+1-i}{\bf 1}$, $10\leqslant i\leqslant 6m+$ 1, in $f_{\alpha_{1}}(1)v^{1}$ are
[…]
and
$(-4+\frac{2}{2m+1})a_{6m+1}-b_{6m}(i=6m+1),\ (-4+\frac{2}{2m+1})a_{0}+2y_{0}(i=0).$
Then (4.6)and (4.7) hold.
The coefficients of $e_{\alpha_{1}}(-1)h_{1}(-1)^{i}h_{2}(-1)^{6m+1-i}{\bf 1}$, $0\leqslant i\leqslant 6m$, in $f_{\alpha_{1}+\alpha_{2}}(1)v^{1}$ are
[…]
This imlpies (4.8).
By (4.10), (4.7), and the fact that $m\geqslant 1$,
$a_{0}=y_{0}=0.$
Then (4.9) follows.
Notice that $(\alpha_{2}|2\alpha_{1}+\alpha_{2})=0$ and $e_{\alpha_{2}}(0)v^{1}=0$. It follows that $f_{\alpha_{2}}(0)v^{1}=0$. Considering the coefficients of $e_{\alpha_{1}}(-1)^{2}h_{1}(-1)^{i}h_{2}(-1)^{6m+1-i}$, $0\leqslant i\leqslant 6m$, we obtain that
[…]
which is (4.10).
The coefficients of $e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)^{2}h_{1}(-1)^{i}h_{2}(-1)^{6-i}{\bf 1}$, $0\leqslant i\leqslant 6m$, in $e_{\alpha+\alpha_{2}}(0)v^{1}$ are
[…]
Then (4.11) follows from (4.2). ∎
By Lemma 4.3, not all $a_{i},x_{j},0\leqslant i\leqslant 6m+1,0\leqslant j\leqslant 6m$ are zero, and by Lemma 4.4, $a_{i}=-x_{i},0\leqslant i\leqslant 6m,\ a_{6m+1}=0$. Then not all $a_{i},0\leqslant i\leqslant 6m+1$ are zero. By (4.10), $a_{0}=0$. Thus we may assume that
$a_{0}=\cdots=a_{r-1}=0,\ a_{r}\neq 0,$
for some $r\geqslant 1$. We have the following lemma.
Lemma 4.5.
For $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, $r=2m.$
Proof.
It can be checked directly that the coefficient of
$e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{r-1}h_{2}(-1)^{6m+1-r}$
in $h_{1}(1)v^{1}$ is
$-5ra_{r}+2r(-3+\frac{2}{2m+1})a_{r}+4x_{r-1}+2y_{r-1}+2b_{r-1}+c_{r-1}+c_{r-2}-z_{r-1}+4l_{r-2}+2(q_{r-1}+q_{r-2}),$ (4.12)
which should be zero. By (4.2) and (4.10),
$x_{r-1}=y_{r-1}=0.$ (4.13)
Then by (4.3) and (4.4), we have
$l_{r-2}=2ra_{r},\ z_{r-1}=-ra_{r}.$
By (4.2) and (4.6),
$b_{r-1}=\frac{-4m}{2m+1}a_{r}.$
By (4.8) and (4.13),
$c_{0}=0,\ c_{j}+c_{j-1}=0,1\leqslant j\leqslant r-1.$
Then we have
$c_{0}=\cdots=c_{r-1}=0.$
By (4.11),
$q_{r-1}+q_{r-2}=ra_{r}.$
Thus (4.12) becomes
$(\frac{4r}{2m+1}-\frac{8m}{2m+1})a_{r}=0.$
This deduces that $r=2m$. ∎
Let $U^{0}$ be the $sl_{3}$-module generated by $v^{1}$. The following lemma is easy to check.
Lemma 4.6.
With respect to the Cartan subalgebra $\mathfrak{h}={\mathbb{C}}h_{1}\oplus{\mathbb{C}}h_{2}$, the weight zero subspace of $U^{0}$ is one-dimensional, which is linearly spanned by $f_{\alpha_{1}}(0)f_{\alpha_{1}+\alpha_{2}}(0)v^{1}$.
Let $W^{1}$ be the subspace of $V^{k}(sl_{3})$ linearly spanned by monomials $z=z^{(+)}z^{(-)}z^{(0)}{\bf 1}$ such that $z\in F^{0}(V^{k}(sl_{3}))$ and $z^{(-)}\neq 1$. Then it can be checked directly that
$f_{\alpha_{1}}(0)f_{\alpha_{1}+\alpha_{2}}(0)v^{1}=\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1}+w^{1},$ (4.14)
for some $w^{1}\in W^{1}+F^{1}(V^{k}(sl_{3}))$. We denote
$v^{3}=f_{\alpha_{1}}(0)f_{\alpha_{1}+\alpha_{2}}(0)v^{1},$
and without loss of generality we may assume that
$a_{2m}=1.$
We have the following lemma.
Lemma 4.7.
$\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1}=h_{1}(-1)^{2m+1}h_{2}(-1)^{2m+1}(h_{1}+h_{2})(-1)^{2m+1}{\bf 1}.$ (4.15)
That is,
$\begin{array}[]{ll}&v^{1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}^{2m+1}(h_{1}+h_{2})(-1)^{2m}\\ \\ &+e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
and
$v^{3}=h_{1}(-1)^{2m+1}h_{2}(-1)^{2m+1}(h_{1}+h_{2})(-1)^{2m+1}{\bf 1}+w^{1},$ (4.16)
where $w^{1}$ is the same as in (4.14).
Proof.
It suffices to prove that
$\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1}=h_{1}(-1)^{2m+1}h_{2}(-1)^{2m+1}(h_{1}+h_{2})(-1)^{2m+1}{\bf 1}.$
Denote
$\pi_{2}=\exp f_{\alpha_{2}}(0)\exp(-e_{\alpha_{2}}(0))\exp f_{\alpha_{2}}(0),$
$\pi_{3}=\exp f_{\alpha_{1}+\alpha_{2}}(0)\exp(-e_{\alpha_{1}+\alpha_{2}}(0))\exp f_{\alpha_{1}+\alpha_{2}}(0).$
Then by Lemma 4.6, there exists $0\neq c_{i}\in{\mathbb{C}}$, $i=2,3$ such that
$\pi_{i}(v^{3})=c_{i}v^{3}.$
On the other hand, it is obvious that
$\pi_{i}(W^{1}+F^{1}(V^{k}(sl_{3})))\subseteq W^{1}+F^{1}(V^{k}(sl_{3})),\ i=2,3.$
So for $i=2,3$, we have
$\begin{array}[]{ll}&\pi_{i}(\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1})\\ \\ =&c_{i}(\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1}).\end{array}$
Notice that
$\begin{array}[]{ll}&\pi_{2}(\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1})\\ \\ =&\sum\limits_{i=2m}^{6m+1}a_{i}(h_{1}+h_{2})(-1)^{i+1}(-h_{2}(-1))^{6m+1-i}h_{1}(-1){\bf 1}.\end{array}$
This means that
$\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1}$
has a factor $(h_{1}+h_{2})(-1)^{2m+1}$. Also
$\begin{array}[]{ll}&\pi_{3}(\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1})\\ \\ =&-\sum\limits_{i=2m}^{6m+1}a_{i}h_{2}(-1)^{i+1}h_{1}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1}.\end{array}$
This implies that $\sum\limits_{i=2m}^{6m+1}a_{i}h_{1}(-1)^{i+1}h_{2}(-1)^{6m+1-i}(h_{1}+h_{2})(-1){\bf 1}$ has a factor $h_{2}(-1)^{2m+1}$. Thus (4.15) holds. ∎
For the singular vector $v^{2}$, we can similarly obtain the following lemma.
Lemma 4.8.
There exists a non-zero number $c$ such that
$\begin{array}[]{ll}&cv^{2}-e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m+1}h_{2}^{2m}(h_{1}+h_{2})(-1)^{2m}\\ \\ &-e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
We are now in a position to give the second main result of this section.
Theorem 4.9.
For $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, $X_{L_{k}(sl_{3})}=\overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}$, where $\theta=\alpha_{1}+\alpha_{2}$.
Proof.
Notice that
$\mathfrak{g}=G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})\cup G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})\cup G\cdot({\mathfrak{h}}^{reg})\cup{\mathcal{N}},$
and
$G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})\cup{\mathcal{N}}\subseteq\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
By Theorem 3.8, the ideal generated by $v^{1}$ and $v^{2}$ is the maximal ideal of $V^{k}(sl_{3})$. Recall that $W^{1}$ is the subspace of $V^{k}(sl_{3})$ linearly spanned by monomials $z=z^{(+)}z^{(-)}z^{(0)}{\bf 1}$ such that $z\in F^{0}(V^{k}(sl_{3}))$ and $z^{(-)}\neq 1$. For $v\in V^{k}(sl_{3})$, denote by $\bar{v}$ the image of $v$ in $R_{V^{k}(sl_{3})}$. Then for any $w\in W^{1}$, the value of $\bar{w}$ at $h_{1}-h_{2}+f_{\theta}$ is zero. Let $h$ be a semisimple regular element in the Cartan subalgebra $\mathfrak{h}={\mathbb{C}}h_{1}+{\mathbb{C}}h_{2}$. Then all the $(h|h_{1})$, $(h|h_{2})$, and $(h|h_{1}+h_{2})$ are non-zeros. Thus by Lemma 4.7,
$h\notin X_{L_{k}(sl_{3})}.$ (4.17)
Then by Lemmas 4.7-4.8, the zero locus of $\overline{{\mathcal{I}}_{k}}\subseteq R_{V^{k}(sl_{3})}$ is exactly $\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$ It follows that
$X_{L_{k}(sl_{3})}=\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
Recall that if $k\in{\mathbb{Z}}_{\geqslant 0}$, $X_{L_{k}(sl_{3})}=\{0\}$ [Zhu96, DM06, Ara12], $X_{L_{k}(sl_{3})}=\overline{{\mathcal{O}}_{min}}$ if $k=-3+\frac{2m+1}{2}$, $m\geqslant 1$ [Ara15a], and $X_{L_{k}(sl_{3})}={\mathcal{N}}$ if $k=-3$ or $k=-3+\frac{p}{q}$, $p,q\geqslant 3$, $(p,q)=1$ [FF92], [Ara15a]. Then we have
Corollary 4.10.
  1. (1)
    For $k\in{\mathbb{C}}$, the associated variety $X_{L_{k}(sl_{3})}$ of $L_{k}(sl_{3})$ is one of the following:
    $\mathfrak{g}^{*},\ \{0\},\ \mathcal{N},\ \overline{{\mathcal{O}}_{min}},\ \overline{G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})},\ \overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
  2. (2)
    Let $h$ be a regular semi-simple vector of $sl_{3}$, then there is no $k\in{\mathbb{C}}$ such that $X_{L_{k}(sl_{3})}=\overline{G\cdot{\mathbb{C}}^{*}h}$.
For $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$, let ${\mathcal{I}}_{k}$ be the maximal ideal of $V^{k}(sl_{3})$, and ${\mathcal{V}}(I_{k})$ be defined as in Section 2. We have the following conjecture.
Conjecture 4.11.
${\mathcal{V}}(I_{k})=\{t\bar{\Lambda}_{1}-\frac{2i}{3}\bar{\Lambda}_{2},t\bar{\Lambda}_{2}-\frac{2i}{3}\bar{\Lambda}_{1},t\bar{\Lambda}_{1}-(t+\frac{2i}{3}+1)\bar{\Lambda}_{2},~{}t\in{\mathbb{C}},i=0,1,\cdots,2m\}.$
Remark 4.12.
If $m=0$, that is $k=-1$, the conjecture is true by Proposition 5.5 of [AP08]. When $m=1$, we could verify the conjecture by computer programming.
5 Varieties of simple affine $W$-algebras $W_{k}(sl_{3},f)$
We first have the following results from [Ara15b] and [Ara13].
Theorem 5.1 ([Ara15b]).
Let $\mathfrak{g}$ be a simple complex Lie algebra, $f$ a regular nilpotent element of $\mathfrak{g}$, and $k=-h^{\vee}+\frac{p}{q}$ a non-degenerate admissible number. Then the simple $W$-algebra $W_{k}(g,f)$ is rational and lisse.
Theorem 5.2 ([Ara13]).
Let $f$ be a minimal nilpotent vector of $sl_{3}$ and $k=-3+\frac{2m+1}{2}$, $m\geqslant 1$. Then the simple $W$-algebra $W_{k}(sl_{3})$ is rational and lisse.
Let $k=-1$ and $f$ a minimal nilpotent element of $sl_{3}$. It was proved in [AKF+18] that $H^{0}_{DS,f}(L_{-1}(sl_{3}))=W_{-1}(sl_{3},f)$, and is isomorphic to the rank one Heisenberg algebra $M(1)$. So its associated variety is one-dimensional.
Theorem 5.3.
Let $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, and $f$ a minimal nilpotent vector of $sl_{3}$, then
  1. (1)
    $X_{W_{k}(sl_{3},f)}=\left\{\left[\begin{array}[]{ccc}a&b&3(\mu^{2}-a^{2})\\ 0&-2a&c\\ 1&0&a\end{array}\right]|~{}\mu,a,b,c\in{\mathbb{C}},\ bc=2(a-\mu)(2a+\mu)^{2}\right\}.$ (5.1)
    In particular,
    $\dim X_{W_{k}(sl_{3},f)}=3.$ (5.2)
  2. (2)
    $W_{k}(sl_{3})$ is not quasi-lisse.
Proof.
By Theorem 2.8, for a minimal vector $f\in sl_{3}$, and $k+3=\frac{2}{2m+1},m\in{\mathbb{Z}}_{\geqslant 0}$,
$H^{0}_{DS,f}(L_{k}(sl_{3}))=W_{k}(sl_{3},f).$ (5.3)
By Theorem 2.9 and (5.3),
$X_{W_{k}(sl_{3}))}=X_{L_{k}(sl_{3})}\cap\mathscr{S}_{f},$
for $k+3=\frac{2}{2m+1}$, $m\in{\mathbb{Z}}_{\geqslant 0}$. We may assume that $f=f_{\theta}$. Then
$\mathscr{S}_{f}=f_{\theta}+\mathfrak{g}^{e_{\theta}},$
and
$\mathfrak{g}^{e_{\theta}}={\mathbb{C}}(\alpha_{1}-\alpha_{2})\oplus{\mathbb{C}}e_{\theta}\oplus{\mathbb{C}}e_{\alpha_{1}}\oplus e_{\alpha_{2}}.$ (5.4)
Then we have
$f_{\theta}+\mathfrak{g}^{e_{\theta}}=\left\{\left[\begin{array}[]{ccc}a&b&d\\ 0&-2a&c\\ 1&0&a\end{array}\right]|~{}a,b,c,d\in{\mathbb{C}}\right\}.$ (5.5)
It is obvious that for $\mu\in{\mathbb{C}}$,
$A_{\mu}=\left[\begin{array}[]{ccc}\mu&0&0\\ 0&-2\mu&0\\ 1&0&\mu\end{array}\right]\in\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
In general,
$B=\left[\begin{array}[]{ccc}a&b&d\\ 0&-2a&c\\ 1&0&a\end{array}\right]\in\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}\cap\mathscr{S}_{f_{\theta}}$
if and only if $B$ is similar to $A_{\mu}$ for some $\mu\in{\mathbb{C}}$. Then we deduce that
$d=3(\mu^{2}-a^{2}),$
and
$bc=2(a-\mu)(\mu+2a)^{2}.$
Thus (5.1) and (5.2) hold.
By Theorem 8.2 of [AFP23a], $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$, are irreducible ordinary modules of $L_{k}(sl_{3})$. This means that $L_{k}(sl_{3})$ has infinitely many irreducible ordinary modules. Then by Theorem 5.4 of [AKR23], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the minimal quantum hamiltonian reductions ${\mathcal{H}}_{j_{t+1,1}\Delta_{t+1,1}}$ and ${\mathcal{H}}_{j_{1,t+1}\Delta_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3},f)$-modules, where
$j_{t+1,1}=\frac{t}{3},\ j_{1,t+1}=\frac{2t}{3},\ \Delta_{t+1,1}=\frac{t^{2}+3t}{3(k+3)}-\frac{2t}{3},\ \Delta_{1,t+1}=\frac{t^{2}+3t}{3(k+3)}-\frac{t}{3}.$
This means that $W_{k}(sl_{3},f)$ has infinitely many irreducible modules. By Theorem 2.3, $W_{k}(sl_{3},f)$ is not quasi-lisse. ∎
We now assume that $f$ is a regular nilpotent vector of $sl_{3}$. We may assume that
$f=\left[\begin{array}[]{ccc}0&0&0\\ 2&0&0\\ 0&2&0\end{array}\right],\ e=\left[\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ 0&0&0\end{array}\right],\ h=\left[\begin{array}[]{ccc}2&0&0\\ 0&0&0\\ 0&0&-2\end{array}\right].$
Then
$\mathfrak{g}^{e}={\mathbb{C}}e\oplus{\mathbb{C}}e_{\theta},$
and
$f+\mathfrak{g}^{e}=\left\{\left[\begin{array}[]{ccc}0&a&b\\ 2&0&a\\ 0&2&0\end{array}\right]|~{}a,b\in{\mathbb{C}}\right\}.$ (5.6)
Denote the associated universal affine $W$-algebras and the simple quotients by $W^{k}(sl_{3})$ and $W_{k}(sl_{3})$, respectively.
Theorem 5.4.
Let $k+3=\frac{2}{2m+1}$ or $\frac{2m+1}{2}$, then
$X_{W_{k}(sl_{3})}=\left\{\left[\begin{array}[]{ccc}0&\frac{3}{4}\mu^{2}&\frac{1}{2}\mu^{3}\\ 2&0&\frac{3}{4}\mu^{2}\\ 0&2&0\end{array}\right]|~{}\mu\in{\mathbb{C}}\right\}.$
In particular, $\dim X_{W_{k}(sl_{3})}=1$. Furthermore, $W_{k}(sl_{3})$ is not quasi-lisse.
Proof.
Since $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, it follows that
$(k\Lambda_{0}+\widehat{\rho}|\alpha)\notin{\mathbb{Z}},\ {\rm for}\ \alpha\in\{-\alpha_{1}+\delta,-\alpha_{2}+\delta,-(\lambda_{1}+\alpha_{2})+\delta,-(\alpha_{1}+\alpha_{2})+2\delta\}.$
Then by Theorem 9.1.4 of [Ara07],
$H^{0}_{DS,f}(L_{k}(sl_{3}))=W_{k}(sl_{3},f).$
Thus
$X_{W_{k}(sl_{3}))}=X_{L_{k}(sl_{3})}\cap\mathscr{S}_{f}.$
By the Feigin-Frenkel Langlands duality [FF92], [ACL19],
$W^{k}(sl_{n})\cong W^{k^{\prime}}(sl_{n}),$
for $k+n=\frac{p}{q}$ and $k^{\prime}+3=\frac{q}{p}$. In particular, for $k+3=\frac{2}{2m+1}$ and $k^{\prime}+3=\frac{2m+1}{2}$,
$W_{k}(sl_{3})\cong W_{k^{\prime}}(sl_{3}).$
Let
$B=\left[\begin{array}[]{ccc}0&a&b\\ 2&0&a\\ 0&2&0\end{array}\right]\in f+\mathfrak{g}^{e}.$
Then
$f_{B}(\lambda)=|\lambda I-B|=\left|\begin{array}[]{ccc}\lambda&-a&-b\\ -2&\lambda&-a\\ 0&-2&\lambda\end{array}\right|=\lambda^{3}-4a\lambda-4b.$
Notice that
$X_{L_{k}(sl_{3})}=\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
So $B\in X_{W_{k}(sl_{3})}$ if and only if $B$ is similar to
$A_{\mu}=\mu(h_{1}-h_{2})+f_{\theta}=\left[\begin{array}[]{ccc}-\mu&0&0\\ 0&2\mu&0\\ 1&0&-\mu\end{array}\right].$
for some $\mu\in{\mathbb{C}}$. Then it is easy to deduce that $B\in X_{W_{k}(sl_{3})}$ if and only if
$a=\frac{3}{4}\mu^{2},\ b=\frac{1}{2}\mu^{3}.$
Thus
$X_{W_{k}(sl_{3})}=\left\{\left[\begin{array}[]{ccc}0&\frac{3}{4}\mu^{2}&\frac{1}{2}\mu^{3}\\ 2&0&\frac{3}{4}\mu^{2}\\ 0&2&0\end{array}\right]|~{}\mu\in{\mathbb{C}}\right\}.$
Let $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$ be the irreducible ordinary modules of $L_{k}(sl_{3})$ as in the proof of Theorem 5.3, then by Theorem 5.4 of [AKR23], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the principal quantum hamiltonian reductions ${\mathcal{W}}_{h_{t+1,1}w_{t+1,1}}$ and ${\mathcal{W}}_{h_{1,t+1}w_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3})$-modules, where
$h_{t+1,1}=\frac{t^{2}+3t}{3(k+3)}=h_{1,t+1},$
and
$w_{t+1,1}=-\frac{2t\sqrt{3}}{3(k+3)^{3}}(\frac{2t}{3}-k-2)(\frac{t}{3}-k-2)=-w_{1,t+1}.$
It follows from Theorem 2.3 that $W_{k}(sl_{3})$ is not quasi-lisse. ∎

### v3

@@
-Cuipo Jiang and Jingtian Song
+Associated varieties of simple affine VOAs $L_{k}(sl_{3})$ and $W$-algebras $W_{k}(sl_{3},f)$
 Abstract
@@
 The associated variety $X_{V}$ of $V$ is the reduced scheme $X_{V}=\operatorname{Specm}(R_{V})$ of $R_{V}$ [Ara12].
-This fundamental invariant of $V$ has been extensively studied in [BFM91, Zhu96, ABD04, Miy04, Ara12, Ara15a, Ara15b, AM18a, AK18, AM17, AJM21, ADF+24, AFK24].
+This fundamental invariant of $V$ has been extensively studied in [BFM91, Zhu96, ABD04, Miy04, Ara12, Ara15a, Ara15b, AM18a, AK18, AM17, AJM21, ADF+25, AFK24].
 Let $\mathfrak{g}$ be a complex simple finite-dimensional Lie algebra and $V^{k}(\mathfrak{g})$ the universal affine vertex algebra at level $k\in{\mathbb{C}}$.
@@
 $k+h^{\vee}=\frac{p}{q},\quad p,q\in{\mathbb{Z}}_{\geqslant 1},\ (p,q)=1,\ p\geqslant\begin{cases}h^{\vee}&\text{if }(r^{\vee},q)=1,\\ h&\text{if }(r^{\vee},q)\neq 1,\end{cases}$
-the variety $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit in $\mathfrak{g}$. In particular, if $k$ is a non-degenerate admissible number, that is,
+the variety $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit in $\mathfrak{g}$ [Ara15a]. In particular, if $k$ is a non-degenerate admissible number, that is,
 $q\geqslant\begin{cases}h&\text{if }(r^{\vee},q)=1,\\ r^{\vee}h^{\vee}&\text{if }(r^{\vee},q)=r^{\vee},\end{cases}$
@@
 If $k=-h^{\vee}$, it follows from [FF92, FG04] that $X_{L_{k}(\mathfrak{g})}$ is also the nilpotent cone.
-It was observed in [AM18a, AM18b, AM17, ADF+24] that there are cases when $L_{k}(\mathfrak{g})$ is non-admissible and $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit, which provide examples for the conjecture in physics [XY21] that, in view of the 4D/2D duality, there should be a large list of non-admissible simple affine vertex algebras whose associated varieties are the closures of some nilpotent orbits.
+It was observed in [AM18a, AM18b, AM17, ADF+25] that there are cases when $L_{k}(\mathfrak{g})$ is non-admissible and $X_{L_{k}(\mathfrak{g})}$ is the closure of some nilpotent orbit, which provide examples for the conjecture in physics [XY21] that, in view of the 4D/2D duality, there should be a large list of non-admissible simple affine vertex algebras whose associated varieties are the closures of some nilpotent orbits.
 There are also cases where $X_{L_{k}(\mathfrak{g})}$ is neither $\mathfrak{g}^{*}$ nor contained in the nilpotent cone ${\mathcal{N}}(\mathfrak{g})$.
 It was proved in [AM17] that $X_{L_{-1}(sl_{n})}$, $n\geqslant 4$, $X_{L_{-m}(sl_{2m})}$, $m\geqslant 2$, $X_{L_{2-r}(so_{2r})}$, $r\in 2{\mathbb{Z}}_{+}+1$, are closures of sheets.
-In [AFK24], it was shown that the associated variety of a simple affine vertex algebras is contained in the closure of the Dixmier sheet when a chiralization of generalized Grothendieck’s simultaneous resolution exists.
+In [AFK24], it was shown that the associated variety of a simple affine vertex algebras is contained in the closure of a Dixmier sheet when a chiralization of generalized Grothendieck’s simultaneous resolution exists.
 The variety $X_{L_{-\frac{5}{2}}(sl_{4})}$ of $L_{-\frac{5}{2}}(sl_{4})$ was given in [Fas22], which turns out to be the closure of a Jordan class of $sl_{4}$.
@@
 In this paper, we will determine $X_{L_{k}(sl_{3})}$ for all non-admissible numbers $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$.
-Then together with the known results $X_{L_{k}(\mathfrak{g})}$ for the critical case and admissible cases given in [FF92, FG04, Ara12], we will obtain a complete characterization of $X_{L_{k}(sl_{3})}$ for all $k\in{\mathbb{C}}$.
+Then together with the results $X_{L_{k}(\mathfrak{g})}$ for the critical case and admissible cases given in [FF92, FG04, Ara12], we will obtain a complete characterization of $X_{L_{k}(sl_{3})}$ for all $k\in{\mathbb{C}}$.
 Another question we are concerned about in this paper is the characterization of the maximal ideal of $V^{k}(sl_{n})$, where $k=-n+\frac{n-1}{q}$, $n\geqslant 3$, $q\geqslant 2$, $(n-1,q)=1$.
-If $q=1$, that is, $k=-1$, the maximal ideal of $V^{-1}(sl_{3})$ is generated by two singular vectors of conformal weight 3 [AP08, AKMF+20], and the maximal ideal of $V^{-1}(sl_{n})$ is generated by one singular vector of conformal weight 2 [AP08, AM17].
-By using the character formulas given in [KT00], the fact that the coefficients as values of Kazhan-Lusztig (or inverse KL) polynomials at 1 in the formulas depend only on the corresponding Coxeter group ( see also Theorem 11 in [Fie06] from the category point of view), and the results for the $k=-1$ case in [KW18, BKK24, AP08, AM17, AKMF+20], we prove our first main result as follows (also see Theorem 3.8).
+If $q=1$, that is, $k=-1$, the maximal ideal of $V^{-1}(sl_{3})$ is generated by two singular vectors of conformal weight 3 [AP08, AKMP+20], and the maximal ideal of $V^{-1}(sl_{n})$ for $n\geqslant 4$ is generated by one singular vector of conformal weight 2 [AP08, AM17].
+By using the character formulas given in [KT00], the fact that the coefficients as values of Kazhan-Lusztig (or inverse KL) polynomials at 1 in the formulas depend only on the corresponding Coxeter group, the category equivalences given in [Fie06], and the results for the $k=-1$ case in [KW18, BKK24, AP08, AM17, AKMP+20], we prove our first main result as follows (also see Theorem 3.11).
 Theorem 1.1.
@@
 $X_{L_{-1}(sl_{3})}=\overline{{\mathbb{S}}_{\mathfrak{l}_{1}}}=\overline{G.{\mathbb{C}}^{*}\lambda}=G.{\mathbb{C}}^{*}\lambda\cup\overline{{\mathcal{O}}_{\min}},$
-where ${\mathfrak{l}}_{1}=\mathfrak{h}\oplus{\mathbb{C}}e_{\theta}\oplus{\mathbb{C}}f_{\theta}$, $\lambda=h_{1}-h_{2}$, and ${\mathcal{O}}_{\min}$ is the minimal non-zero nilpotent orbit of $sl_{3}({\mathbb{C}})$.
+where ${\mathfrak{l}}_{1}=\mathfrak{h}\oplus{\mathbb{C}}e_{\theta}\oplus{\mathbb{C}}f_{\theta}$, $\lambda=h_{1}-h_{2}$, and ${\mathcal{O}}_{\min}$ is the minimal non-zero nilpotent orbit of $sl_{3}$.
 In particular, $X_{L_{-1}(sl_{3})}$ is irreducible and $\dim X_{L_{-1}(sl_{3})}=5$.
+Remark 1.3.
+The result of Theorem 1.2 is also covered in [AFK24] in a different and more conceptual way.
 For $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$, we have our third main result as follows (see also Theorem 4.9).
@@
 We will study the associated varieties of the simple vertex operator algebras $V^{k}(sl_{n})$ for $n\geqslant 4$ and general non-admissible numbers $k$ in subsequent papers.
-Representations of $L_{k}(sl_{3})$ have been extensively studied in [AP08, Ada16, AFR17, Cre18, FLLZ18, AM21, AK21, AKR21, FKR21, FR22, KRW22, AFP23a, CRR23, ACG24], etc.
-We here have the following conjecture for $L_{k}(sl_{3})$ (also see Conjecture 4.11).
+Representations of $L_{k}(sl_{3})$ have been extensively studied in [AP08, Ada16, AFR17, Cre18, FLLZ18, AM21, AK21, AKR21, FKR21, FR22, KRW22, AMP23, CRR23, ACG24, AKR24], etc.
+We have the following conjecture for $L_{k}(sl_{3})$ (also see Conjecture 4.11).
 Conjecture 1.5.
@@
 If $k=-3+\frac{2m+1}{2}$, $m\geqslant 1$, it was shown in [Ara13] that $X_{W_{k}(sl_{3},f)}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational.
-If $k=-1$, it has been proved in [AKF+18] that $W_{-1}(sl_{3},f)$ is isomorphic to the rank one Heisenberg vertex operator algebra.
+If $k=-1$, it has been proved in [AKMP+18] that $W_{-1}(sl_{3},f)$ is isomorphic to the rank one Heisenberg vertex operator algebra.
 So $X_{W_{-1}(sl_{3},f)}$ is one-dimensional.
@@
 Let $f$ be a regular nilpotent element of $\mathfrak{g}=sl_{3}$.
-If $k=-3+\frac{p}{q}$, $p,q\geqslant 3$, $(p,q)=1$, by [Ara15b], $X_{W_{k}(sl_{3})}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational.
+If $k=-3+\frac{p}{q}$, $p,q\geqslant 3$, $(p,q)=1$, by [Ara15b], $X_{W_{k}(sl_{3},f)}=\{0\}$ and $W_{k}(sl_{3},f)$ is rational.
 Our fifth main result is given as follows (also see Theorem 5.4).
@@
 Let $k+3=\frac{2}{2m+1}$ or $\frac{2m+1}{2}$, $m\geqslant 1$, and $f$ a regular nilpotent element of $sl_{3}$. Then
-$X_{W_{k}(sl_{3})}=\left\{\left[\begin{array}[]{ccc}0&\frac{3}{4}\mu^{2}&\frac{1}{2}\mu^{3}\\ 2&0&\frac{3}{4}\mu^{2}\\ 0&2&0\end{array}\right]|~{}\mu\in{\mathbb{C}}\right\}.$
-In particular, $\dim X_{W_{k}(sl_{3})}=1$. Furthermore, $W_{k}(sl_{3},f)$ is not quasi-lisse.
+$X_{W_{k}(sl_{3},f)}=\left\{\left[\begin{array}[]{ccc}0&\frac{3}{4}\mu^{2}&\frac{1}{2}\mu^{3}\\ 2&0&\frac{3}{4}\mu^{2}\\ 0&2&0\end{array}\right]|~{}\mu\in{\mathbb{C}}\right\}.$
+In particular, $\dim X_{W_{k}(sl_{3},f)}=1$. Furthermore, $W_{k}(sl_{3},f)$ is not quasi-lisse.
 The rest of this paper is organized as follows.
@@
 In Section 3, we study singular vectors and maximal ideals of $V^{k}(sl_{n})$ for $k=-n+\frac{n-1}{q}$ with $n\geqslant 3$, $q\geqslant 1$, $(n-1,q)=1$.
-Section 4 is dedicated to determine the associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$.
+Section 4 is dedicated to determining the associated varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$.
 In Section 5, we obtain the associated varieties of the simple affine $W$-algebras $W_{k}(sl_{3},f)$.
 Acknowledgements
-We appreciate Tomoyuki Arakawa, Dražen Adamović, and Anne Moreau for valuable discussions and comments.
-In particular, the first author is grateful to Dražen Adamović for showing her the very helpful result of Theorem 8.2 in his paper [AFP23b] joint with Pierluigi Möseneder Frajria and Paolo Papi, and to Tomoyuki Arakawara for telling her the papers [Fie06] and [BKK24], and for the discussions.
+We appreciate Tomoyuki Arakawa, Dražen Adamović, and Anne Moreau for invaluable discussions and comments.
+In particular, the first author is grateful to Dražen Adamović for showing her the very helpful result of Theorem 8.2 in his paper [AMP23] joint with Pierluigi Möseneder Frajria and Paolo Papi, and to Tomoyuki Arakawara for telling her the papers [Fie06] and [BKK24], and for the discussions.
+Results in this paper were presented in part in conferences “Representation Theory XVIII” in Dubrovnik, 2023; “Vertex Algebras , Infinite Dimensional Lie Algebras, and Related Topics” in Rome, 2023; “RIMS Joint Research Activity: Vertex Algebras and Related Topics ” in Kyoto, 2024; and “TYMRC-Workshop on Quantum Algebras and Representations” in Shenzhen, 2024.
+We thank the organizers of these conferences.
+Finally, we are very grateful to the referees for invaluable comments and suggestions.
 2 Preliminaries
@@
 Furthermore, if $S$ is a finite set, then $V$ is called finitely strongly generated by $S$.
-For a vertex algebra $V$, define $C_{2}(V)=\operatorname{span}_{{\mathbb{C}}}\{a_{-2}b\mid a,b\in V\}$ and
+For a vertex algebra $V$, let $C_{2}(V)=\operatorname{span}_{{\mathbb{C}}}\{a_{-2}b\mid a,b\in V\}$ and
 $R_{V}=V/C_{2}(V).$
@@
 Definition 2.2 ([AK18]).
-Let $V$ be a vertex algebra. Then $V$ is called quasi-lisse if $X_{V}$ has finitely many symplectic leaves.
+Let $V$ be a vertex algebra. $V$ is called quasi-lisse if $X_{V}$ has finitely many symplectic leaves.
 Theorem 2.3 ([AK18]).
-Let $V$ be a finitely strongly generated vertex algebra. If $V$ is quasi-lisse, then $V$ has finitely many ordinary modules.
+Let $V$ be a finitely strongly generated vertex algebra. If $V$ is quasi-lisse, then $V$ has finitely many irreducible ordinary modules.
 2.2 Affine vertex algebras
@@
 $V^{k}(\mathfrak{g})\cong U(\mathfrak{g}\otimes t^{-1}{\mathbb{C}}[t^{-1}])=U(t^{-1}\mathfrak{g}[t^{-1}]).$ (2.1)
-The space $V^{k}(\mathfrak{g})$ is naturally graded,
+The space $V^{k}(\mathfrak{g})$ is naturally graded as follows:
 $V^{k}(\mathfrak{g})=\bigoplus_{n\in{\mathbb{Z}}_{\geqslant 0}}V^{k}(\mathfrak{g})_{n},$
@@
 For $\lambda\in\mathfrak{h}^{*}$, let $L(\lambda)$ be the irreducible module of $V^{k}(\mathfrak{g})$ induced from the simple highest weight $\mathfrak{g}$-module $L(\bar{\lambda})$.
-Then by [Zhu96], $L(\lambda)$ is a module of $L_{k}(\mathfrak{g})$ if and only if $I_{k}(L(\bar{\lambda}))=0$.
+Then by [Zhu96], $L(\lambda)$ is a module of $L_{k}(\mathfrak{g})$ if and only if $I_{k}$ acts on $L(\bar{\lambda})$ as zero.
 We have the following result from [Ara16].
@@
 $F^{1}V=C_{2}(V),\qquad R_{V}=V/C_{2}(V)=F^{0}V/F^{1}V\subset{\rm gr}^{F}V.$
-It is easily seen that
+It is easy to see that
 $F^{1}V^{k}(\mathfrak{g})=C_{2}(V^{k}(\mathfrak{g}))=t^{-2}\mathfrak{g}[t^{-1}]V^{k}(\mathfrak{g}).$
@@
 Identifying $\mathfrak{g}^{*}$ with $\mathfrak{g}$ via the bilinear form (|)($\cdot$|$\cdot$)( ⋅ | ⋅ ), one may view $X_{V}$ as a subvariety of $\mathfrak{g}$.
-2.6 Sheet and Jordan classes
+2.6 Sheets and Jordan classes
 In this subsection, we recall Jordan classes and sheets of a semisimple Lie algebras following [TY05].
@@
 $W^{k}(\mathfrak{g},f)=H^{0}_{DS,f}(V^{k}(\mathfrak{g})).$
-Here $H^{\bullet}_{DS,f}(M)$ denotes the BRST cohomology of the generalized quantized Drinfeld-Sokolov reduction associated with $f\in\mathcal{N}(\mathfrak{g})$ with coefficients in a $V^{k}(\mathfrak{g})$-module $M$ [KRW03].
+Here $H^{\bullet}_{DS,f}(M)$ denotes the BRST cohomology of the generalized quantized Drinfeld-Sokolov reduction associated to $f\in\mathcal{N}(\mathfrak{g})$ with coefficients in a $V^{k}(\mathfrak{g})$-module $M$ [KRW03].
 Recall that we have a natural isomorphism $R_{W^{k}(\mathfrak{g},f)}\cong{\mathbb{C}}[\mathscr{S}_{f}]$ of Poisson algebras [DSK06, Ara15a], so that
@@
   1. (1)
-    $H_{DS,f_{\theta}}^{i}(M)=0$ for all $i\neq 0$, $M\in{\mathcal{O}}_{k}$. In particular, the functor
+    $H_{DS,f_{\theta}}^{i}(M)=0$ for all $i\neq 0$, and $M\in{\mathcal{O}}_{k}$. In particular, the functor
     $\mathcal{O}_{k}\longrightarrow W^{k}(\mathfrak{g},f_{\theta})\operatorname{-Mod},\quad M\longmapsto H^{0}_{DS,f_{\theta}}(M),$
@@
   1. (1)
-    $H_{DS,f}^{i}(M)=0$ for all $i\neq 0$, $M\in\operatorname{KL}_{k}$. In particular, the functor
+    $H_{DS,f}^{i}(M)=0$ for all $i\neq 0$, and $M\in\operatorname{KL}_{k}$. In particular, the functor
     $\operatorname{KL}_{k}\longrightarrow W^{k}(\mathfrak{g},f)\operatorname{-Mod},\quad M\longmapsto H^{0}_{DS,f}(M),$
@@
 For $k\in{\mathbb{C}}$, $V^{k}(\mathfrak{g})$ is not simple if and only if
-$r(k+h^{\vee})\in{\mathbb{Q}}_{\geqslant 0}\backslash\left\{\frac{1}{m}\mid m\in{\mathbb{Z}}_{\geqslant 1}\right\},$ (3.1)
-where $r$ is the lacing number of $\mathfrak{g}$.
+$r^{\vee}(k+h^{\vee})\in{\mathbb{Q}}_{\geqslant 0}\backslash\left\{\frac{1}{m}\mid m\in{\mathbb{Z}}_{\geqslant 1}\right\},$ (3.1)
+where $r^{\vee}$ is the lacing number of $\mathfrak{g}$.
 By Theorem 3.1, $V^{k}(sl_{n})$ is not simple if and only if $k=-n+\frac{p}{q}$, where $p\in{\mathbb{Z}}_{\geqslant 2}$, $q\in{\mathbb{Z}}_{\geqslant 1}$, and $(p,q)=1$.
 In this section, we will prove that, for $k=-3+\frac{2}{2m+1}$, $m\in{\mathbb{Z}}_{\geqslant 1}$, the maximal ideal of $V^{k}(sl_{3})$ is generated by two singular vectors of conformal weight $3(2m+1)$, and for $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$, the maximal ideal of $V^{k}(sl_{n})$ is generated by a singular vector of conformal weight $2q$.
-We reach our results by using the Kashiwara-Tanisaki characters [KT00] (see also [Fie06]), the character formulas for the $k=-1$ case given in [KW18, BKK24], and the results established in [AP08, AKMF+20, AM17].
+We reach our results by using the Kashiwara-Tanisaki characters [KT00], category equivalences established in [Fie06], the character formulas for the $k=-1$ case given in [KW18, BKK24], and the results established in [AP08, AKMP+20, AM17].
 3.1 Kashiwara-Tanisaki Characters
@@
 Denote the Bruhat ordering of $\widehat{W}(\lambda)$ by $\geqslant_{\lambda}$.
-For $y,w\in\widehat{W}(\lambda)$, denote by $P^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the corresponding Kazhdan-Lusztig polynomial [KL79], and by $Q^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the inverse Kazhdan-Lusztig polynomial defined by
+For $y,w\in\widehat{W}(\lambda)$, denote by $P^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the associated Kazhdan-Lusztig polynomial [KL79], and by $Q^{\lambda}_{y,w}(q)\in{\mathbb{Z}}[q]$ the inverse Kazhdan-Lusztig polynomial defined by
 $\sum\limits_{x\leqslant_{\lambda}y\leqslant_{\lambda}z}(-1)^{l_{\lambda}(y)-l_{\lambda}(x)}Q^{\lambda}_{x,y}(q)P^{\lambda}_{y,z}(q)=\delta_{x,z},$
@@
 simple module) of $\widehat{\mathfrak{g}}$ with highest weight $\lambda$.
-We have the following result from [KT00].
+The following lemma comes from [KK79], [DGK82], [Kum87], [KW89].
+Lemma 3.2.
+Let $\lambda\in{\mathcal{C}}^{+}$, and $w\in\widehat{W}(\lambda)$.
+If $L(\mu)$ is a subquotient of $M(w\circ\lambda)$, then there exists $y\in\widehat{W}(\lambda)$ such that $\mu=yw\circ\lambda$, and $w\leqslant_{\lambda}yw$.
+We have the following character formulas from [KT00].
 Theorem 3.3.
@@
     ${\rm ch}(L(w\circ\lambda))=\sum\limits_{w\geqslant_{\lambda}y\in\widehat{W}(\lambda)}(-1)^{l_{\lambda}(w)-l_{\lambda}(y)}P^{\lambda}_{y,w}(1){\rm ch}(M(y\circ\lambda)).$
+We have the following lemma from [KT98], [Fie06].
+Lemma 3.4.
+Let $\lambda\in{\cal C}^{+}$ or ${\cal C}^{-}$, and $w,w^{\prime}\in\widehat{W}(\lambda)$ be such that $w^{\prime}\cdot\lambda\leqslant w\cdot\lambda$. Then
+${\rm dim}{\rm Hom}_{\mathbb{C}}(M(w^{\prime}\cdot\lambda),M(w\cdot\lambda))=1.$ (3.2)
+The following theorem comes from [Fie06].
+Theorem 3.5.
+Let ${\mathfrak{h}}\subset{\mathfrak{b}}\subset{\mathfrak{g}}$ and ${\mathfrak{h}}^{\prime}\subset{\mathfrak{b}}^{\prime}\subset{\mathfrak{g}}^{\prime}$ be symmetrizable Kac-Moody Lie algebras with Cartan and Borel subalgebras, respectively.
+Let $\Lambda\in\mathfrak{h}^{*}/\sim$ and $\Lambda^{\prime}\in(\mathfrak{h}^{\prime})^{*}/\sim^{\prime}$ be two equivalent classes outside the critical hyperplanes and let ${\cal O}_{\Lambda}$ and ${\cal O}^{\prime}_{\Lambda^{\prime}}$ be the corresponding indecomposable blocks, where $\sim$ and $\sim^{\prime}$ are the usual relation in $\mathfrak{h}^{*}$ and $(\mathfrak{h}^{\prime})^{*}$, respectively (see [KK79]).
+Suppose the following conditions hold:
+  1. (1)
+    There exist $\lambda\in\Lambda$ and $\lambda^{\prime}\in\Lambda^{\prime}$ which are either both dominant or antidominant;
+  2. (2)
+    $(W(\lambda),S(\lambda))\cong(W(\lambda^{\prime}),S(\lambda^{\prime}))$ are isomorphic as Coxeter groups;
+  3. (3)
+    $W_{0}(\lambda)\cong W_{0}(\lambda^{\prime})$ under the same isomorphism and both sets are finite.
+Then there exists an equivalence of categories
+${\cal O}_{\Lambda}\cong{\cal O^{\prime}}_{\Lambda^{\prime}}.$
+Remark 3.6.
+As an application of Theorem 3.5, the author in [Fie06] proved Theorem 3.3 for symmetrizable Kac-Moody Lie algbera $\mathfrak{g}$, and regular $\Lambda\in{\mathfrak{h}}^{*}/\sim$ satisfying the condition in Theroem 3.5 and that $(W(\Lambda),S(\Lambda))$ is finite or affine type, and dominant or antidominant $\lambda\in\Lambda$.
 […]
@@
 as Coxeter groups by $\sigma:(\widehat{W},S)\to(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$ such that
-$\sigma s_{\alpha_{0}}=s_{\beta_{0}},\sigma(s_{\alpha_{i}})=s_{\alpha_{i}},\ i=1,2,\cdots,n-1.$
+$\sigma s_{\alpha_{0}}=s_{\beta_{0}},\sigma s_{\alpha_{i}}=s_{\alpha_{i}},\ i=1,2,\cdots,n-1.$
+In particular, $\sigma(W_{0}(-\Lambda_{0}))=W_{0}(k\Lambda_{0})$.
 Proof.
@@
 $t_{\gamma}(\lambda)=\lambda+(\lambda|\delta)\gamma-((\lambda|\gamma)+\frac{1}{2}(\lambda|\delta)(\gamma|\gamma))\delta,$
-for $\lambda\in\widehat{\mathfrak{h}}^{*}$. Recall from [Kac90], for any $x\in\widehat{W}$, there exists $y\in W$ and $\gamma\in Q$ such that $x=yt_{\gamma}$. We have the following lemma.
+for $\lambda\in\widehat{\mathfrak{h}}^{*}$.
+Recall from [Kac90], for any $x\in\widehat{W}$, there exists $y\in W$ and $\gamma\in Q$ such that $x=yt_{\gamma}$. We have the following lemma.
 Lemma 3.8.
@@
 Theorem 3.9 ([KW18, BKK24]).
-Let $n\geqslant 3$, and let $L(\Lambda)$ be an irreducible level $k=-1$ $\widehat{sl}_{n}$-module with highest weight
+Let $n\geqslant 3$, and let $L(\Lambda)$ be an irreducible $\widehat{sl}_{n}$-module at level $k=-1$ with highest weight
 $\Lambda=-(1+s)\Lambda_{0}+s\Lambda_{n-1}\ ({\rm resp.}\ \Lambda=-(1+s)\Lambda_{0}+s\Lambda_{1}),\ s\in{\mathbb{Z}}_{\geqslant 0}.$
 Then the character of $L(\Lambda)$ is given by the following formula:
-$\widehat{R}{\rm ch}L(\Lambda)=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1}(resp.\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}({\Lambda+\widehat{\rho}})-\widehat{\rho}}.$
+$\widehat{R}{\rm ch}L(\Lambda)=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1}(resp.\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}({\Lambda+\widehat{\rho}})-\widehat{\rho}},$
 where
@@
 In particular, if $\Lambda=-\Lambda_{0}$,
-$\widehat{R}{\rm ch}L(-\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
-We also need the following results from [KK79], [DGK82], [Kum87], [KW89].
-Lemma 3.6.
-Let $\lambda\in{\mathcal{C}}^{+}$, and $w\in\widehat{W}(\lambda)$. If $L(\mu)$ is a subquotient of $M(w\circ\lambda)$, then there exists $y\in\widehat{W}(\lambda)$ such that $\mu=yw\circ\lambda$.
-Recall the following results from [AP08], [AKMF+20], and [AM17].
+$\widehat{R}{\rm ch}L(-\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$ (3.9)
+Recall the following results from [AP08], [AKMP+20], and [AM17].
 Theorem 3.10.
@@
     [AP08] Let
-    $u=e_{\theta}(-1)e_{\alpha_{2}+\cdots\alpha_{n-2}}(-1){\bf 1}-e_{\alpha_{1}+\cdots\alpha_{n-2}}(-1)e_{\alpha_{2}+\cdots\alpha_{n-1}}(-1){\bf 1}.$
+    $u=e_{\theta}(-1)e_{\alpha_{2}+\cdots+\alpha_{n-2}}(-1){\bf 1}-e_{\alpha_{1}+\cdots+\alpha_{n-2}}(-1)e_{\alpha_{2}+\cdots+\alpha_{n-1}}(-1){\bf 1}.$
     Then $u$ is a singular vector of $V^{-1}(sl_{n})$, $n\geqslant 4$.
   3. (3)
-    [AKMF+20] The ideal of $V^{-1}(sl_{3})$ generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$.
+    [AKMP+20] The ideal of $V^{-1}(sl_{3})$ generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$.
   4. (4)
@@
 Proof.
-Let $\mathfrak{g}=sl_{n}$, $n\geqslant 3$.
+Let $\mathfrak{g}=sl_{n}$, $n\geqslant 3$. It is obvious that $-\Lambda_{0},k\Lambda_{0}\in{\mathcal{C}}^{+}$, for $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$.
 Notice that if $q=1$, then $k=-1$ and $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))=(\widehat{W},S)$.
-So $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))\cong(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2,(q,n-1)=1$.
-By Lemma 3.3, and the definitions of $P^{\lambda}_{y,w}(q)$ and $Q^{\lambda}_{y,w}(q)$ [KL79], we have for $y,w\in\widehat{W}$,
+So by Lemma 3.7, $(\widehat{W},S)\cong(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$ through $\sigma$, for $k=-n+\frac{n-1}{q}$, $q\geqslant 2,(q,n-1)=1$.
+In particular,
+$\sigma(s_{\alpha_{0}})=s_{\beta_{0}},\ \sigma(s_{\alpha_{i}})=s_{\alpha_{i}},\ i=1,2,\cdots,n-1.$ (3.11)
+Then for $w,y\in\widehat{W}$,
 $w\geqslant_{-\Lambda_{0}}y\ {\rm if\ and\ if}\ \sigma(w)\geqslant_{k\Lambda_{0}}\sigma(y),$ (3.12)
-$l_{-\Lambda_{0}}(y)=l_{k\Lambda_{0}}\sigma(y),\ l_{-\Lambda_{0}}(w)=l_{k\Lambda_{0}}\sigma(w),$ (3.10)
 and
+$l_{-\Lambda_{0}}(y)=l_{k\Lambda_{0}}\sigma(y),\ l_{-\Lambda_{0}}(w)=l_{k\Lambda_{0}}\sigma(w).$ (3.13)
+Let $w\in\widehat{W}(-\Lambda_{0})$ be such that $w$ is the longest element of $w\widehat{W}_{0}(-\Lambda_{0})$.
+Then by (3.11) and (3.13), $\sigma(w)$ is the longest element of $\sigma(w)\widehat{W}_{0}(k\Lambda_{0})$.
+By Theorem 3.3,
+${\rm ch}(L(w\circ(-\Lambda_{0})))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1){\rm ch}(M(y\circ(-\Lambda_{0}))),$ (3.14)
+and
+${\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{\sigma(w)\leqslant_{k\Lambda_{0}}\sigma(y)\in\widehat{W}(k\Lambda_{0})}(-1)^{l_{k\Lambda_{0}}(\sigma(y))-l_{k\Lambda_{0}}(\sigma(w))}Q^{k\Lambda_{0}}_{\sigma(w),\sigma(y)}(1){\rm ch}(M(\sigma(y)\circ k\Lambda_{0})).$ (3.15)
+Since $(\widehat{W}(-\Lambda_{0}),S(-\Lambda_{0}))\cong(\widehat{W}(k\Lambda_{0}),S(k\Lambda_{0}))$ via the automorphism $\sigma$, by the definition of $P^{\lambda}_{y,w}(q)$ and $Q^{\lambda}_{y,w}(q)$ [KL79], for $y,w\in\widehat{W}(-\Lambda_{0})$,
 $P^{-\Lambda_{0}}_{y,w}(1)=P^{k\Lambda_{0}}_{\sigma(y),\sigma(w)}(1),\ Q^{-\Lambda_{0}}_{y,w}(1)=Q^{k\Lambda_{0}}_{\sigma(y),\sigma(w)}(1).$ (3.16)
-It is obvious that $-\Lambda_{0},k\Lambda_{0}\in{\mathcal{C}}^{+}$, for $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$.
-Notice that for any $w\in\widehat{W}(-\Lambda_{0})$ which is the longest element of $w\widehat{W}_{0}(-\Lambda_{0})$, $\sigma(w)$ is the longest element of $\sigma(w)\widehat{W}_{0}(k\Lambda_{0})$.
-Then by Theorem 3.2 and (3.9)-(3.11) ( see also Theorem 11 of [Fie06] from the category point of view),
-${\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1){\rm ch}(M(\sigma(y)\circ(k\Lambda_{0}))),$ (3.12)
-which is equivalent to
-$\widehat{R}{\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1)e^{\sigma(y)\circ k\Lambda_{0}}.$ (3.13)
-Then by Lemma 3.3, Lemma 3.4, and Theorem 3.5,
-$\widehat{R}{\rm ch}L(k\Lambda_{0})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{q\gamma}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
-Furthermore, (3.12) together with Lemma 3.6 means that $L(y\circ(-\Lambda_{0}))$ for some $y\in\widehat{W}(-\Lambda_{0})$ is isomorphic to a subquotient of $M(w\circ(-\Lambda_{0}))$ if and only if $L(\sigma(y)\circ k\Lambda_{0})$ is a subquotient of $M(\sigma(w)\circ k\Lambda_{0})$.
+Then together with (3.12)-(3.13), we have
+${\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1){\rm ch}(M(\sigma(y)\circ k\Lambda_{0})).$ (3.17)
+The equation systems (3.14) and (3.17) also implies that for $w,y\in\widehat{W}$,
+$[M(w\circ(-\Lambda_{0})):L(y\circ(-\Lambda_{0}))]=[M(\sigma(w)\circ k\Lambda_{0}):L(\sigma(y)\circ k\Lambda_{0})],$ (3.18)
+where $[M:L(\mu)]$ denotes the multiplicity of the irreducible subquotient $L(\mu)$ in $M$, for a module $M$ in category ${\cal O}$.
+By the character formulas given in [KK79] for generalized Verma modules of affine Lie algebras induced from integrable modules of the associated simple Lie algebras, one can easily deduce that
+$[V(w\circ(-\Lambda_{0})):L(y\circ(-\Lambda_{0}))]=[V(\sigma(w)\circ k\Lambda_{0}):L(\sigma(y)\circ k\Lambda_{0})],$ (3.19)
+where $V(w\circ(-\Lambda_{0}))$ is the generalized Verma module of the affine Lie algebra $\widehat{sl_{n}}$ induced from the finite-dimensional $sl_{n}$-module $L(\overline{w\circ(-\Lambda_{0}}))$, and $V(\sigma(w)\circ(k\Lambda_{0}))$ is the generalized Verma module induced from $L(\overline{\sigma(w)\circ k\Lambda_{0}})$.
+Notice that (3.17) can also be described as
+$\widehat{R}{\rm ch}(L(\sigma(w)\circ k\Lambda_{0}))=\sum\limits_{w\leqslant_{-\Lambda_{0}}y\in\widehat{W}(-\Lambda_{0})}(-1)^{l_{-\Lambda_{0}}(y)-l_{-\Lambda_{0}}(w)}Q^{-\Lambda_{0}}_{w,y}(1)e^{\sigma(y)\circ k\Lambda_{0}},$ (3.20)
+where
+$\widehat{R}=\prod_{\alpha\in\widehat{\Delta}_{+}}(1-e^{-\alpha})^{{\rm dim}{\widehat{\mathfrak{g}}}_{\alpha}},\ \mathfrak{g}=sl_{n}.$
+By (3.9),
+$\widehat{R}{\rm ch}L_{-1}(sl_{n}))=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{\gamma}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
+By Lemma 3.7 and Lemma 3.8, for $w\in W$, $\gamma\in Q$,
+$\sigma(wt_{\gamma})=wt_{q\gamma}.$
+Then we have
+$\widehat{R}{\rm ch}L_{k}(sl_{n})=\sum\limits_{w\in W}\sum\limits_{\tiny{\begin{gathered}\gamma\in Q\\ (\gamma|\bar{\Lambda}_{n-1})\geqslant 0,(\gamma|\bar{\Lambda}_{1}))\geqslant 0\end{gathered}}}(-1)^{l(w)}e^{wt_{q\gamma}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho}}.$
+(3.10) is proved.
+Set
+$\Lambda^{(-1)}=\widehat{W}(-\Lambda_{0})\circ(-\Lambda_{0})=\{w\circ(-\Lambda_{0})|w\in\widehat{W}(-\Lambda_{0})\}$
+and
+$\Lambda^{(k)}=\widehat{W}(k\Lambda_{0})\circ k\Lambda_{0}=\{w\circ(k\Lambda_{0})|w\in\widehat{W}(k\Lambda_{0})\}.$
+Let ${\cal O}_{\Lambda^{(-1)}}$ and ${\cal O}_{\Lambda^{(k)}}$ be the corresponding indecomposable blocks, respectively. Then by Theorem 3.5 and Lemma 3.7,
+${\cal O}_{\Lambda^{(-1)}}\cong{\cal O}_{\Lambda^{(k)}}.$ (3.21)
+We need to explain the equivalence (3.21) briefly.
+Notice that both $-\Lambda_{0}$ and $k\Lambda_{0}$ are dominant.
+Then for $w\in\widehat{W}(-\Lambda_{0})$, $L(w\circ(-\Lambda_{0}))$ (resp.
+$L(\sigma(w)\circ k\Lambda_{0})$) has a projective cover $P(w\circ(-\Lambda_{0}))$ (resp.
+$P(\sigma(w)\circ k\Lambda)$) in ${\cal O}_{\Lambda^{(-1)}}$ (resp.
+${\cal O}_{\Lambda^{(k)}}$).
+Set
+${\cal P}_{\Lambda^{(-1)}}=\{P(w\circ(-\Lambda_{0}))|w\in\widehat{W}(-\Lambda_{0})\}$
+and
+${\cal P}_{\Lambda^{(k)}}=\{P(\sigma(w)\circ k\Lambda_{0})|w\in\widehat{W}(-\Lambda_{0})\}.$
+${\cal P}_{\Lambda^{(-1)}}$ (resp.
+${\cal P}_{\Lambda^{(k)}}$) is a faithful set of small projectives in the sense of [Mit72].
+View ${\cal P}_{\Lambda^{(-1)}}$ (resp.
+${\cal P}_{\Lambda^{(k)}}$) as a full subcategory of ${\cal O}_{\Lambda^{(-1)}}$(resp.
+${\cal O}_{\Lambda^{(k)}}$).
+Let ${\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(-1)}}^{\rm opp}}$ (resp.
+${\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(k)}}^{\rm opp}}$) be the category of additive functors ${\cal P}_{\Lambda^{(-1)}}\rightarrow{{\mathbb{C}}}-{\rm mod}$ (resp.
+${\cal P}_{\Lambda^{(k)}}\rightarrow{{\mathbb{C}}}-{\rm mod}$).
+By Theorem 3.1 of [Mit72],
+${\cal O}_{\Lambda^{(-1)}}\rightarrow{\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(-1)}}^{\rm opp}}\ \ ({\rm resp.}\ {\cal O}_{\Lambda^{(k)}}\rightarrow{\mathbb{C}}-{\rm mod}^{{\cal P}_{\Lambda^{(k)}}^{\rm opp}})$
+$M\rightarrow{\rm Hom}(\cdot,\ M)\ \ \ ({\rm resp.}\ \tilde{M}\rightarrow{\rm Hom}(\cdot,\ \tilde{M})).$
+is an equivalence of ${{\mathbb{C}}}$-categories.
+Theorem 3.5 (i.e., Theorem 11 of [Fie06]) tells us that the Coxeter group automorphism $\sigma$ determines an equivalence between the categories ${\cal O}_{\Lambda^{(-1)}}$ and ${\cal O}_{\Lambda^{(k)}}$ via the above ${\mathbb{C}}$-category equivalences (c.f.
+[Fie06] for details).
+Let $F$ be the covariant functor from ${\cal O}_{\Lambda^{(-1)}}$ to ${\cal O}_{\Lambda^{(k)}}$.
+Then in particular, $F$ maps the class of Verma module $M(w\circ(-\Lambda_{0}))$ to the class of Verma module $M(\sigma(w)\circ k\Lambda_{0})$, the class of simple module $L(w\circ(-\Lambda_{0}))$ to the class of simple module $L(\sigma(w)\circ k\Lambda_{0})$, and it is easy to see that $F$ maps the class of $V(w\circ(-\Lambda_{0}))$ to the class of $V(\sigma(w)\circ k\Lambda_{0})$, for $w\in\widehat{W}(-\Lambda_{0})$.
 Case 1   $n=3$, $k=-3+\frac{2}{2m+1}$, $m\geqslant 1$.
-Let $u^{1}$ and $u^{2}$ be the singular vectors given in (1) of Theorem 3.7.
-The weights of $u^{1}$ and $u^{2}$ with respect to $\widehat{\mathfrak{h}}$ are $-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2}$ and $-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}$, respectively.
-It is easily checked that
-$s_{\alpha_{2}}\circ(t_{\alpha_{1}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2},$ (3.14)
+Let $u^{1}$ and $u^{2}$ be the singular vectors given in (1) of Theorem 3.10 with highest weights $-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2}$ and $-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}$, respectively.
+It is easy to check that
+$s_{\alpha_{2}}\circ t_{\alpha_{1}}\circ(-\Lambda_{0})=s_{\alpha_{2}}\circ(t_{\alpha_{1}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+2\alpha_{1}+\alpha_{2},$ (3.22)
 and
-$s_{\alpha_{1}}\circ(t_{\alpha_{2}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}.$ (3.15)
-Thus $V^{k}(sl_{3})$ has two singular vectors $v^{1}$ and $v^{2}$ of wights $s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$ and $s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$, respectively.
+$s_{\alpha_{1}}\circ t_{\alpha_{2}}\circ(-\Lambda_{0})=s_{\alpha_{1}}\circ(t_{\alpha_{2}}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=-\Lambda_{0}-3\delta+\alpha_{1}+2\alpha_{2}.$ (3.23)
+Then by (3.18) and Lemma 3.4, $V^{k}(sl_{3})$ has two singular vectors $v^{1}$ and $v^{2}$ of weights $s_{\alpha_{2}}\circ t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0}$ and $s_{\alpha_{1}}\circ t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0}$, respectively.
 It is easy to check that
-$s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2},$ (3.16)
+$s_{\alpha_{2}}\circ t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0}=s_{\alpha_{2}}\circ(t_{(2m+1)\alpha_{1}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+2\alpha_{1}+\alpha_{2},$ (3.24)
 and
-$s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2},$ (3.17)
-By (3) of Theorem 3.7, the ideal generated by $u^{1}$ and $u^{2}$ is the maximal ideal of $V^{-1}(sl_{3})$.
-It follows that the ideal generated by $v^{1}$ and $v^{2}$ is the maximal ideal of $V^{k}(sl_{3})$.
-We complete the proof of (1).
+$s_{\alpha_{1}}\circ t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0}=s_{\alpha_{1}}\circ(t_{(2m+1)\alpha_{2}}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-3(2m+1)\delta+\alpha_{1}+2\alpha_{2}.$ (3.25)
+Since
+$[V^{-1}(sl_{3}):L(s_{\alpha_{1}}t_{\alpha_{2}}\circ(-\Lambda_{0}))]=[V^{-1}(sl_{3}):L(s_{\alpha_{2}}t_{\alpha_{1}}\circ(-\Lambda_{0}))]=1,$
+by (3.19),
+$[V^{k}(sl_{3}):L(s_{\alpha_{1}}t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0})]=[V^{k}(sl_{3}):L(s_{\alpha_{2}}t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0})]=1.$ (3.26)
+Let $\tilde{I}$ be the ideal of $V^{k}(sl_{3})$ generated by $v^{1}$ and $v^{2}$. Since $I$ is the maximal ideal of $V^{-1}(sl_{3})$, we have the following exact sequence:
+$0\rightarrow I\rightarrow V^{-1}(sl_{3})\rightarrow L_{-1}(sl_{3})\rightarrow 0$
+Then we have the exact sequence:
+$0\rightarrow F(I)\rightarrow V^{k}(sl_{3})\rightarrow L_{k}(sl_{3})\rightarrow 0,$
+since $F(V^{-1}(sl_{3}))=V^{k}(sl_{3})$, $F(L_{-1}(sl_{3}))=L_{k}(sl_{3})$.
+On the other hand, since $I$ is a quotient of $M(s_{\alpha_{1}}\circ t_{\alpha_{2}}\circ(-\Lambda_{0}))\oplus M(s_{\alpha_{2}}\circ t_{\alpha_{1}}\circ(-\Lambda_{0}))$, it follows that $F(I)$ is a quoitent of $M(s_{\alpha_{1}}\circ t_{(2m+1)\alpha_{2}}\circ k\Lambda_{0})\oplus M(s_{\alpha_{2}}\circ t_{(2m+1)\alpha_{1}}\circ k\Lambda_{0})$.
+This together with (3.26) means that $F(I)$ must be the ideal $\tilde{I}$ generated by $v^{1}$ and $v^{2}$.
+Since $F(I)$ is maximal, it follows that $\tilde{I}$ is maximal.
 Case 2  $n\geqslant 4$, $k=-n+\frac{n-1}{q}$, $q\geqslant 2$, $(q,n-1)=1$.
@@
 The weight of $u$ with respect to the Cartan subalgebra $\widehat{\mathfrak{h}}$ is $-\Lambda_{0}-2\delta+\theta+\beta$, where $\beta=\alpha_{2}+\cdots+\alpha_{n-2}$.
-It is easily checked that
+It is easy to check that
 $-\Lambda_{0}-2\delta+\theta+\beta=(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}})\circ(t_{-\beta}(-\Lambda_{0}+\widehat{\rho})-\widehat{\rho}).$
-Then $V^{k}(sl_{n})$ has a singular vector $w^{1}$ of weight $(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}})\circ(t_{-q\beta}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})$. It can be checked directly that
+Then by (3.18) and Lemma 3.4, $V^{k}(sl_{n})$ has a singular vector $w^{1}$ of weight $(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-q\beta})\circ k\Lambda_{0}$. One can easily check that
 $(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}})\circ(t_{-q\beta}(k\Lambda_{0}+\widehat{\rho})-\widehat{\rho})=k\Lambda_{0}-2q\delta+\theta+\beta.$
-By (4) of Theorem 3.7, the ideal generated by $u$ is the maximal ideal of $V^{-1}(sl_{n})$. It follows that the ideal generated by $w^{1}$ is the maximal ideal of $V^{k}(sl_{n})$. ∎
-Remark 3.9.
-The method here can be used to the subregular case for $\mathfrak{g}$ of other types.
+By (4) of Theorem 3.10, the ideal $J$ generated by $u$ is the maximal ideal of $V^{-1}(sl_{n})$. Then we have the exact sequence:
+$0\rightarrow J\rightarrow V^{-1}(sl_{n})\rightarrow L_{-1}(sl_{n})\rightarrow 0$ (3.27)
+By the fact that $F(V^{-1}(sl_{n}))=V^{k}(sl_{n})$, and $F(L_{-1}(sl_{n}))=L_{k}(sl_{n})$, we have
+$0\rightarrow F(J)\rightarrow V^{k}(sl_{n})\rightarrow L_{k}(sl_{n})\rightarrow 0$ (3.28)
+Notice that $J$ is a quotient of $M(s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-\beta})\circ(-\Lambda_{0}))$, it follows that $F(J)$ is a quotient of $M((s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-q\beta})\circ k\Lambda_{0})$.
+Then $F(J)$ is generated by $w^{1}$, since
+$[V^{-1}(sl_{n}):L((s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-\beta})\circ(-\Lambda_{0}))]=[V^{k}(sl_{n}):L((s_{\theta}s_{\alpha_{1}}s_{\alpha_{n-1}}t_{-q\beta})\circ k\Lambda_{0})]=1.$
+By (3.28), $F(J)$ is maximal. We complete the proof of (2). We remark that (3.28) can also be deduced from (3.27) and (3.19) without using the category equivalence (3.21). ∎
 4 Varieties of $L_{k}(sl_{3})$ for $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$
@@
 $\begin{array}[]{ll}u^{1}=&[-e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)+e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{2}(-1)+e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{2}}(-1)]{\bf 1},\\ \\ u^{2}=&[e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}+e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)\\ &-2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-2)-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}}(-1)]{\bf 1}.\end{array}$
-are singular vectors of $V^{-1}(sl_{3})$. Then we have
+are singular vectors of $V^{-1}(sl_{3})$. Then we have the following lemma.
 Lemma 4.1.
 There exist $x,y\in{\mathcal{I}}_{-1}$ such that
-$\begin{array}[]{ll}x\equiv&-e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)h_{2}(-1){\bf 1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1){\bf 1}\\ \\ &-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}+\alpha_{2}}(-1){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)(2h_{1}+h_{2})(-1){\bf 1}\\ \\ &+2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{2}}(-1){\bf 1}({\rm mod}~{}F^{1}(V^{-1}(sl_{3})))\\ \\ y\equiv&h_{1}(-1)h_{2}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{3}}(-1){\bf 1}\\ \\ &-3e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1)f_{\alpha_{2}}(-1){\bf 1}-e_{\alpha_{1}}(-1)f_{\alpha_{1}}(-1)h_{1}(-1){\bf 1}\\ \\ &+e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{2}}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}\\ \\ &-e_{\alpha_{2}}(-1)f_{\alpha_{2}}(-1)h_{2}(-1){\bf 1}\ ({\rm mod}~{}F^{1}(V^{-1}(sl_{3}))).\end{array}$
+$\begin{array}[]{ll}x\equiv&-e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)h_{2}(-1){\bf 1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1){\bf 1}\\ \\ &-e_{\alpha_{1}+\alpha_{2}}(-1)^{2}f_{\alpha_{1}+\alpha_{2}}(-1){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)(2h_{1}+h_{2})(-1){\bf 1}\\ \\ &+2e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{2}}(-1){\bf 1}({\rm mod}~{}F^{1}(V^{-1}(sl_{3})))\\ \\ y\equiv&h_{1}(-1)h_{2}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}+e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{2}}(-1){\bf 1}\\ \\ &-3e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}}(-1)f_{\alpha_{2}}(-1){\bf 1}-e_{\alpha_{1}}(-1)f_{\alpha_{1}}(-1)h_{1}(-1){\bf 1}\\ \\ &+e_{\alpha_{1}+\alpha_{2}}(-1)f_{\alpha_{1}+\alpha_{2}}(-1)(h_{1}(-1)+h_{2}(-1)){\bf 1}\\ \\ &-e_{\alpha_{2}}(-1)f_{\alpha_{2}}(-1)h_{2}(-1){\bf 1}\ ({\rm mod}~{}F^{1}(V^{-1}(sl_{3}))).\end{array}$
 Proof.
@@
 $\begin{array}[]{ll}p_{1}=&-e_{\alpha_{1}}^{2}e_{\alpha_{2}}+e_{\alpha_{1}}e_{\alpha_{1}+\alpha_{2}}h_{2}+e_{\alpha_{1}+\alpha_{2}}^{2}f_{\alpha_{2}},\\ \\ p_{2}=&e_{\alpha_{1}}e_{\alpha_{2}}^{2}+e_{\alpha_{2}}e_{\alpha_{1}+\alpha_{2}}h_{1}-e_{\alpha_{1}+\alpha_{2}}^{2}f_{\alpha_{1}},\\ \\ p_{3}=&-e_{\alpha_{1}+\alpha_{2}}h_{1}h_{2}-e_{\alpha_{1}}e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{1}}-e_{\alpha_{1}+\alpha_{2}}^{2}f_{\alpha_{1}+\alpha_{2}}+e_{\alpha_{1}}e_{\alpha_{2}}(2h_{1}+h_{2})+2e_{\alpha_{2}}e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{2}},\\ \\ p_{4}=&h_{1}h_{2}(h_{1}+h_{2})+e_{\alpha_{1}}e_{\alpha_{2}}f_{\alpha_{1}+\alpha_{3}}-3e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{1}}f_{\alpha_{2}}-e_{\alpha_{1}}f_{\alpha_{1}}h_{1},\\ \\ &+e_{\alpha_{1}+\alpha_{2}}f_{\alpha_{1}+\alpha_{2}}(h_{1}+h_{2})-e_{\alpha_{2}}f_{\alpha_{2}}h_{2}.\end{array}$
-Then by Lemma 4.1, $p_{i}\in\overline{{\mathcal{I}}_{-1}}$, $1\leqslant i\leqslant 4$.
+Then by Lemma 4.1, $p_{i}\in\overline{{\mathcal{I}}_{-1}}$, $1\leqslant i\leqslant 4$. We have the following result which is also covered in [AFK24] by different method.
 Theorem 4.2.
@@
 $p_{1}(f_{\alpha_{1}}+f_{\alpha_{2}})=-1,\ p_{3}(t\lambda+f_{\theta})=-t^{2}.$
-This implies that $G\cdot(f_{\alpha_{1}}+f_{\alpha_{2}})$ and $G\cdot({\mathbb{C}}^{*}\lambda+f_{\theta})$ could not belong in $X_{L_{-1}(sl_{3})}$.
+This implies that $G.(f_{\alpha_{1}}+f_{\alpha_{2}})$ and $G.({\mathbb{C}}^{*}\lambda+f_{\theta})$ could not belong in $X_{L_{-1}(sl_{3})}$.
 Furthermore, for regular semisimple vector $h$, all $(h_{1}|h),\ (h|h_{2}),(h|h_{1}+h_{2})$ are non-zero.
@@
 […]
-This imlpies (4.8).
+This implies (4.8).
 By (4.10), (4.7), and the fact that $m\geqslant 1$,
@@
 That is,
-$\begin{array}[]{ll}&v^{1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}^{2m+1}(h_{1}+h_{2})(-1)^{2m}\\ \\ &+e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
+$\begin{array}[]{ll}&v^{1}-e_{\alpha_{1}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}^{2m+1}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\\ \\ &+e_{\alpha_{1}}(-1)^{2}e_{\alpha_{2}}(-1)h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
 and
@@
 There exists a non-zero number $c$ such that
-$\begin{array}[]{ll}&cv^{2}-e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m+1}h_{2}^{2m}(h_{1}+h_{2})(-1)^{2m}\\ \\ &-e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
+$\begin{array}[]{ll}&cv^{2}-e_{\alpha_{2}}(-1)e_{\alpha_{1}+\alpha_{2}}(-1)h_{1}(-1)^{2m+1}h_{2}^{2m}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\\ \\ &-e_{\alpha_{1}}(-1)e_{\alpha_{2}}(-1)^{2}h_{1}(-1)^{2m}h_{2}(-1)^{2m}(h_{1}+h_{2})(-1)^{2m}{\bf 1}\in W^{1}+F^{1}(V^{k}(sl_{3})).\end{array}$
 We are now in a position to give the second main result of this section.
@@
 Notice that
-$\mathfrak{g}=G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})\cup G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})\cup G\cdot({\mathfrak{h}}^{reg})\cup{\mathcal{N}},$
+$\mathfrak{g}=G.{\mathbb{C}}^{*}(h_{1}-h_{2})\cup G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})\cup G\cdot({\mathfrak{h}}^{reg})\cup{\mathcal{N}},$
 and
-$G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})\cup{\mathcal{N}}\subseteq\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
+$G.{\mathbb{C}}^{*}(h_{1}-h_{2})\cup{\mathcal{N}}\subseteq\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
 By Theorem 3.11, the ideal generated by $v^{1}$ and $v^{2}$ is the maximal ideal of $V^{k}(sl_{3})$.
@@
 $h\notin X_{L_{k}(sl_{3})}.$ (4.17)
-Then by Lemmas 4.7-4.8, the zero locus of $\overline{{\mathcal{I}}_{k}}\subseteq R_{V^{k}(sl_{3})}$ is exactly $\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$ It follows that
-$X_{L_{k}(sl_{3})}=\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
+Then by Lemmas 4.7-4.8, the zero locus of $\overline{{\mathcal{I}}_{k}}\subseteq R_{V^{k}(sl_{3})}$ is exactly $\overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$ It follows that
+$X_{L_{k}(sl_{3})}=\overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
 Recall that if $k\in{\mathbb{Z}}_{\geqslant 0}$, $X_{L_{k}(sl_{3})}=\{0\}$ [Zhu96, DM06, Ara12], $X_{L_{k}(sl_{3})}=\overline{{\mathcal{O}}_{min}}$ if $k=-3+\frac{2m+1}{2}$, $m\geqslant 1$ [Ara15a], and $X_{L_{k}(sl_{3})}={\mathcal{N}}$ if $k=-3$ or $k=-3+\frac{p}{q}$, $p,q\geqslant 3$, $(p,q)=1$ [FF92], [Ara15a].
@@
     For $k\in{\mathbb{C}}$, the associated variety $X_{L_{k}(sl_{3})}$ of $L_{k}(sl_{3})$ is one of the following:
-    $\mathfrak{g}^{*},\ \{0\},\ \mathcal{N},\ \overline{{\mathcal{O}}_{min}},\ \overline{G\cdot{\mathbb{C}}^{*}(h_{1}-h_{2})},\ \overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
+    $\mathfrak{g}^{*},\ \{0\},\ \mathcal{N},\ \overline{{\mathcal{O}}_{min}},\ \overline{G.{\mathbb{C}}^{*}(h_{1}-h_{2})},\ \overline{G.({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
   2. (2)
-    Let $h$ be a regular semi-simple vector of $sl_{3}$, then there is no $k\in{\mathbb{C}}$ such that $X_{L_{k}(sl_{3})}=\overline{G\cdot{\mathbb{C}}^{*}h}$.
+    Let $h$ be a regular semi-simple vector of $sl_{3}$, then there is no $k\in{\mathbb{C}}$ such that $X_{L_{k}(sl_{3})}=\overline{G.{\mathbb{C}}^{*}h}$.
 For $k=-3+\frac{2}{2m+1}$, $m\geqslant 0$, let ${\mathcal{I}}_{k}$ be the maximal ideal of $V^{k}(sl_{3})$, and ${\mathcal{V}}(I_{k})$ be defined as in Section 2. We have the following conjecture.
 Conjecture 4.11.
-${\mathcal{V}}(I_{k})=\{t\bar{\Lambda}_{1}-\frac{2i}{3}\bar{\Lambda}_{2},t\bar{\Lambda}_{2}-\frac{2i}{3}\bar{\Lambda}_{1},t\bar{\Lambda}_{1}-(t+\frac{2i}{3}+1)\bar{\Lambda}_{2},~{}t\in{\mathbb{C}},i=0,1,\cdots,2m\}.$
+${\mathcal{V}}(I_{k})=\{t\bar{\Lambda}_{1}-\frac{2i}{2m+1}\bar{\Lambda}_{2},t\bar{\Lambda}_{2}-\frac{2i}{2m+1}\bar{\Lambda}_{1},t\bar{\Lambda}_{1}-(t+\frac{2i}{2m+1}+1)\bar{\Lambda}_{2},~{}t\in{\mathbb{C}},i=0,1,\cdots,2m\}.$
 Remark 4.12.
-If $m=0$, that is $k=-1$, the conjecture is true by Proposition 5.5 of [AP08]. When $m=1$, we could verify the conjecture by computer programming.
+[…]
 5 Varieties of simple affine $W$-algebras $W_{k}(sl_{3},f)$
@@
 Let $\mathfrak{g}$ be a simple complex Lie algebra, $f$ a regular nilpotent element of $\mathfrak{g}$, and $k=-h^{\vee}+\frac{p}{q}$ a non-degenerate admissible number.
-Then the simple $W$-algebra $W_{k}(g,f)$ is rational and lisse.
+Then the simple $W$-algebra $W_{k}(\mathfrak{g},f)$ is rational and lisse.
 Theorem 5.2 ([Ara13]).
@@
 Let $k=-1$ and $f$ a minimal nilpotent element of $sl_{3}$.
-It was proved in [AKF+18] that $H^{0}_{DS,f}(L_{-1}(sl_{3}))=W_{-1}(sl_{3},f)$, and is isomorphic to the rank one Heisenberg algebra $M(1)$.
+It was proved in [AKMP+18] that $H^{0}_{DS,f}(L_{-1}(sl_{3}))=W_{-1}(sl_{3},f)$, and is isomorphic to the rank one Heisenberg algebra $M(1)$.
 So its associated variety is one-dimensional.
@@
   2. (2)
-    $W_{k}(sl_{3})$ is not quasi-lisse.
+    $W_{k}(sl_{3},f)$ is not quasi-lisse.
 Proof.
@@
 By Theorem 2.9 and (5.3),
-$X_{W_{k}(sl_{3}))}=X_{L_{k}(sl_{3})}\cap\mathscr{S}_{f},$
+$X_{W_{k}(sl_{3},f)}=X_{L_{k}(sl_{3})}\cap\mathscr{S}_{f},$
 for $k+3=\frac{2}{2m+1}$, $m\in{\mathbb{Z}}_{\geqslant 0}$. We may assume that $f=f_{\theta}$. Then
@@
 $A_{\mu}=\left[\begin{array}[]{ccc}\mu&0&0\\ 0&-2\mu&0\\ 1&0&\mu\end{array}\right]\in\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}.$
-In general,
+In general, non-semisimple
 $B=\left[\begin{array}[]{ccc}a&b&d\\ 0&-2a&c\\ 1&0&a\end{array}\right]\in\overline{G\cdot({\mathbb{C}}^{*}(h_{1}-h_{2})+f_{\theta})}\cap\mathscr{S}_{f_{\theta}}$
@@
 Thus (5.1) and (5.2) hold.
-By Theorem 8.2 of [AFP23a], $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$, are irreducible ordinary modules of $L_{k}(sl_{3})$.
+By Theorem 8.2 of [AMP23], $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$, are irreducible ordinary modules of $L_{k}(sl_{3})$.
 This means that $L_{k}(sl_{3})$ has infinitely many irreducible ordinary modules.
-Then by Theorem 5.4 of [AKR23], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the minimal quantum hamiltonian reductions ${\mathcal{H}}_{j_{t+1,1}\Delta_{t+1,1}}$ and ${\mathcal{H}}_{j_{1,t+1}\Delta_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3},f)$-modules, where
+Then by Theorem 5.4 of [AKR24], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the minimal quantum hamiltonian reductions ${\mathcal{H}}_{j_{t+1,1}\Delta_{t+1,1}}$ and ${\mathcal{H}}_{j_{1,t+1}\Delta_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3},f)$-modules, where
 $j_{t+1,1}=\frac{t}{3},\ j_{1,t+1}=\frac{2t}{3},\ \Delta_{t+1,1}=\frac{t^{2}+3t}{3(k+3)}-\frac{2t}{3},\ \Delta_{1,t+1}=\frac{t^{2}+3t}{3(k+3)}-\frac{t}{3}.$
@@
 Then by Theorem 9.1.4 of [Ara07],
-$H^{0}_{DS,f}(L_{k}(sl_{3}))=W_{k}(sl_{3},f).$
+$H^{0}_{DS,f}(L_{k}(sl_{3}))=W_{k}(sl_{3}).$
 Thus
@@
 So $B\in X_{W_{k}(sl_{3})}$ if and only if $B$ is similar to
-$A_{\mu}=\mu(h_{1}-h_{2})+f_{\theta}=\left[\begin{array}[]{ccc}-\mu&0&0\\ 0&2\mu&0\\ 1&0&-\mu\end{array}\right].$
+$A_{\mu}=\mu(h_{1}-h_{2})+f_{\theta}=\left[\begin{array}[]{ccc}-\mu&0&0\\ 0&2\mu&0\\ 1&0&-\mu\end{array}\right]$
 for some $\mu\in{\mathbb{C}}$. Then it is easy to deduce that $B\in X_{W_{k}(sl_{3})}$ if and only if
@@
 $X_{W_{k}(sl_{3})}=\left\{\left[\begin{array}[]{ccc}0&\frac{3}{4}\mu^{2}&\frac{1}{2}\mu^{3}\\ 2&0&\frac{3}{4}\mu^{2}\\ 0&2&0\end{array}\right]|~{}\mu\in{\mathbb{C}}\right\}.$
-Let $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$ be the irreducible ordinary modules of $L_{k}(sl_{3})$ as in the proof of Theorem 5.3, then by Theorem 5.4 of [AKR23], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the principal quantum hamiltonian reductions ${\mathcal{W}}_{h_{t+1,1}w_{t+1,1}}$ and ${\mathcal{W}}_{h_{1,t+1}w_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3})$-modules, where
+Let $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$, $t\in{\mathbb{Z}}_{\geqslant 0}$ be the irreducible ordinary modules of $L_{k}(sl_{3})$ as in the proof of Theorem 5.3, then by Theorem 5.4 of [AKR24], for $t\in{\mathbb{Z}}_{\geqslant 0}$, the principal quantum hamiltonian reductions ${\mathcal{W}}_{h_{t+1,1}w_{t+1,1}}$ and ${\mathcal{W}}_{h_{1,t+1}w_{1,t+1}}$ of $L(t\bar{\Lambda}_{1})$ and $L(t\bar{\Lambda}_{2})$ are irreducible ordinary $W_{k}(sl_{3})$-modules, where
 $h_{t+1,1}=\frac{t^{2}+3t}{3(k+3)}=h_{1,t+1},$
@@
 It follows from Theorem 2.3 that $W_{k}(sl_{3})$ is not quasi-lisse. ∎
+6 Declarations
+  1. (1)
+    This work is supported by Natural Science Foundation of China (Grant number: 12171312);
+  2. (2)
+    The authors have no relevant financial or non-financial interests to disclose;
+  3. (3)
+    All datasets supporting the analysis and conclusions of the paper are publicly available at the time of publication.